### v1

# Weighted Fisher divergence for high-dimensional Gaussian variational inference

Abstract
Bayesian inference has many advantages for complex models. However, standard Monte Carlo methods for summarizing the posterior can be computationally demanding, and it is attractive to consider optimization-based variational approximations. Our work considers Gaussian approximations with sparse precision matrices which are tractable to optimize in high-dimensional problems. Although the optimal Gaussian approximation is usually defined as the one closest to the target posterior in Kullback-Leibler divergence, it is useful to consider other divergences when the Gaussian assumption is crude, in order to capture important features of the posterior for a given application. Our work studies the weighted Fisher divergence, which focuses on gradient differences between the target posterior and its approximation, with the Fisher and score-based divergences being special cases. We make three main contributions. First, we compare approximations for weighted Fisher divergences under mean-field assumptions for both Gaussian and non-Gaussian targets with Kullback-Leibler approximations. Second, we go beyond mean-field and consider approximations with sparse precision matrices reflecting posterior conditional independence structure for hierarchical models. Using stochastic gradient descent to enforce sparsity, we develop two approaches to minimize the weighted Fisher divergence, based on the reparametrization trick and a batch approximation of the objective. Finally, we examine the performance of our methods for examples involving logistic regression, generalized linear mixed models and stochastic volatility models.
Fisher divergence,
Score-based divergence,
Stochastic gradient descent,
Gaussian variational approximation,
keywords:
[…]
, and
1 Introduction
Bayesian inference is a powerful tool for quantifying uncertainty, but it is demanding to implement for two reasons. Firstly, specifying a full probabilistic model for all unknowns and observables requires careful thought, and the components of the model need to be checked against the data. Secondly, Bayesian computations are difficult, requiring approximation of high-dimensional integrals. For many Bayesian models, exact posterior inference is infeasible, and a variety of numerical methods for summarizing the posterior are used in practice, such as Markov chain Monte Carlo (MCMC) and variational inference (VI). MCMC is asymptotically unbiased, which means that we can estimate posterior quantities as precisely as we wish with a large enough number of iterations. While MCMC is often treated as the gold standard for posterior estimation, its computational cost can be prohibitively high for large datasets or complex models (Robert and Casella, 2004; Maclaurin and Adams, 2015). On the other hand, VI reformulates posterior computation into an optimization problem by minimizing a divergence between the true posterior and a simpler variational distribution. This enables faster and more scalable inference, leveraging advances in optimization algorithms (Blei et al., 2017). As a result, VI is increasingly popular for its computational efficiency in large-scale problems.
The performance of VI is largely determined by the family of variational approximations chosen, optimization technique, and divergence characterizing discrepancy between the true posterior and variational density. Much of the VI literature has focused on improving expressiveness of the variational family and enhancing optimization methods, often using Kullback-Leibler divergence (KLD) as the measure of approximation quality. To better capture the dependence structure among variables, which can be especially strong in hierarchical models, structured variational approximations that mimic the true dependency structure can be employed (Hoffman and Blei, 2015; Tan and Nott, 2018; Durante and Rigon, 2019; Tan, 2021). More recently, flow-based methods which transform an initial simple distribution into more flexible forms through a series of invertible transformations have been introduced (Rezende et al., 2014; Dinh et al., 2017; Agrawal and Domke, 2024). These approaches allow VI to capture highly complex posterior distributions, significantly enhancing the flexibility of the inference.
Despite the popularity of KLD, studying alternatives can be important, particularly when using simple variational families which may be employed for tractability in high-dimensional problems. These approximations may not be capable of matching the target posterior closely, and choosing an appropriate divergence can help to capture the most important features of the posterior for a given application. A family of divergences including KLD as a special case is the Rényi’s $\alpha$ family (Li and Turner, 2016), with parameter $\alpha$ that can be adjusted to give Hellinger distance ($\alpha=0.5$), $\chi^{2}$-divergence ($\alpha=2$) and KLD ($\alpha=1$). The value of $\alpha$ can be chosen to balance between mode-seeking and mass-covering behavior, but when $\alpha\neq 1$, the most practical methods for optimizing the variational Rényi bound use biased stochastic gradients. Stein divergence has also emerged as a powerful objective for VI. Ranganath et al. (2016) introduced operator variational inference, a minimax approach that optimizes Stein discrepancies by constructing variational objectives based on Stein operators. Liu and Wang (2016) developed Stein variational gradient descent, which uses kernelized Stein discrepancies to iteratively transform particles toward the posterior. In this article, we explore use of the weighted Fisher divergence in Gaussian VI, focusing on the Fisher and score-based divergences as special cases. The definitions and motivations for studying these divergences are presented below.
1.1 Weighted Fisher divergence
Let $p(y|\theta)$ be the likelihood of observed data $y$, where $\theta\in\mathbb{R}^{d}$ is an unknown model parameter. Consider Bayesian inference with a prior density $p(\theta)$. In classical variational inference (Ormerod and Wand, 2010; Blei et al., 2017), the true posterior $p(\theta|y)=p(y|\theta)p(\theta)/p(y)$ is approximated with a more tractable density $q(\theta)$ by minimizing the KLD between them, where
$\text{\rm KL}(q\|p)=\int q(\theta)\log\frac{q(\theta)}{p(\theta|y)}d\theta.$
Let ${\text{E}}_{q}$ denote expectation with respect to $q(\theta)$. As $\log p(y)=\text{\rm KL}(q\|p)+\mathcal{L}$, where $\mathcal{L}={\text{E}}_{q}\{\log p(y,\theta)-\log q(\theta)\}$, minimizing the KLD is equivalent to maximizing an evidence lower bound $\mathcal{L}$ on $\log p(y)$, which does not depend on the normalizing constant of the true posterior.
Score matching (Hyvärinen, 2005) focuses instead on closeness between gradients of the log densities with respect to the variable $\theta$, although the score function refers conventionally to gradient of the log-likelihood with respect to the parameter. A form of such discrepancy is the weighted Fisher divergence (Barp et al., 2019), defined as
[…]
$={\text{E}}_{q}\{\|\nabla_{\theta}\log q(\theta)-\nabla_{\theta}\log p(\theta|y)\|_{M}^{2}\},$
where $\|\cdot\|_{M}$ is the $M$-weighted vector norm defined as $\|z\|_{M}=\sqrt{z^{\top}Mz}$ and $M$ is a positive semi-definite matrix. Like KLD, $S_{M}(q\|p)$ is asymmetric, non-negative, and vanishes when $q(\theta)=p(\theta|y)$. Let $h(\theta)=p(y|\theta)p(\theta)$. Then $\nabla_{\theta}\log p(\theta|y)=\nabla_{\theta}\log h(\theta)$, which is independent of the unknown normalizing constant $p(y)$. Similarly, if the variational density contains an unknown normalizing constant, this does not need to be computed to evaluate the weighted Fisher divergence. Unlike the evidence lower bound, the weighted Fisher divergence provides a direct measure of the distance between the true posterior and variational density.
When $M$ is the identity matrix $I$, $S_{I}(q\|p)$ is known as the Fisher divergence (FD, Hyvärinen, 2005), denoted hereafter as $F(q\|p)$. When $q(\theta)$ is ${\text{N}}(\mu,\Sigma)$ and $M$ is its covariance matrix $\Sigma$, $S_{\Sigma}(q\|p)$ is referred to as score-based divergence (SD) in Cai et al. (2024), denoted as $S(q\|p)$ henceforth. Cai et al. (2024) derived closed-form updates for Gaussian variational parameters in a batch and match (BaM) algorithm based on score-based divergence, and showed that $S(q\|p)$ is affine invariant while $F(q\|p)$ is not. This means that $S(\tilde{q}\|\tilde{p})=S(q\|p)$ if $\tilde{p}$ and $\tilde{q}$ denote the densities of $p$ and $q$ respectively after applying an affine transformation of $\theta$.
In recent years, there is increasing interest in use of the weighted Fisher divergence in VI. Huggins et al. (2020) showed that the Fisher divergence defined in terms of the generalized $\ell_{p}$ norm is an upper bound to the $p$-Wasserstein distance, and its optimization ensures closeness of the variational density to the true posterior in terms of important point estimates and uncertainties. Yang et al. (2019) derived an iteratively reweighted least squares algorithm for minimizing the Fisher divergence in exponential family based variational approximations. Elkhalil et al. (2021) employed the factorizable polynomial exponential family as variational approximation in their Fisher autoencoder framework, which has performance competitive with existing ones. Modi et al. (2023) developed Gaussian score matching variational inference with closed form updates, by minimizing the KLD between a target and Gaussian variational density subject to a matching score function constraint. For implicit variational families structured hierarchically, Yu and Zhang (2023) used the Fisher divergence to reformulate the optimization objective into a minimax problem. Cai et al. (2024) proposed a variational family built on orthogonal function expansions, and transformed the optimization objective into a minimum eigenvalue problem using Fisher divergence.
Our contributions in this article are threefold. First, we study behavior of the weighted Fisher divergence in mean-field Gaussian VI for both Gaussian and non-Gaussian targets, showing its tendency to underestimate the posterior variance more severely than KLD. Second, we develop Gaussian VI for high-dimensional hierarchical models where posterior conditional independence structure is captured via a sparse precision matrix. Sparsity is enforced by computing updates using stochastic gradient descent (SGD) and two approaches are proposed for minimizing the weighted Fisher divergence. The first relies on the reparametrization trick (Kingma and Welling, 2014), while the second relies on a biased estimate of the objective based on a batch of samples at each iteration (Elkhalil et al., 2021; Cai et al., 2024). Third, we study the variance of unbiased gradient estimates computed using the reparametrization trick and limiting behavior of the batch approximated weighted Fisher divergence under the mean-field assumption.
This article is organized as follows. We study the quality of posterior mode and variance approximations for Gaussian and non-Gaussian targets in Sections 2 and 3 respectively, when using the weighted Fisher divergence in VI. Section 4 introduces Gaussian VI for hierarchical models by using a sparse precision matrix to capture posterior conditional independence. Two SGD approaches for minimizing the weighted Fisher divergence are proposed in Sections 5 and 6, based respectively on the reparametrization trick and batch approximation. Experimental results are discussed in Section 7 with applications to logistic regression, generalized linear mixed models (GLMMs) and stochastic volatility models. Section 8 concludes the paper with a discussion of the findings and future work.
2 Ordering of divergences for
Gaussian target
Accurate estimation of the posterior variance is important in VI, as it affects uncertainty quantification in Bayesian inference. Here, we establish an ordering of the weighted Fisher and KL divergences according to the estimated posterior variance when the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$ with a non-diagonal precision matrix $\Lambda$. All divergences considered can recover the true mean $\nu$ and precision $\Lambda$ when the variational family is also Gaussian with a full covariance matrix. However, the computational cost of optimizing a full-rank Gaussian variational approximation can be prohibitive for high-dimensional models. A widely used alternative is the mean-field Gaussian variational approximation, $q(\theta)={\text{N}}(\mu,\Sigma)$, with a diagonal covariance matrix $\Sigma$. The mean-field assumption simplifies the optimization but tends to underestimate the true posterior variance under the KLD (Blei et al., 2017; Tan and Nott, 2018; Giordano et al., 2018). Here, we examine the severity of posterior variance underestimation under the weighted Fisher divergence compared to KLD.
Our results in this section generalize similar results in Margossian et al. (2024) from score-based divergences to the more general class of weighted Fisher divergences. For KLD under the mean-field assumption, Margossian et al. (2024) showed that the posterior mean can be recovered ($\mu=\nu$) and the optimal variance parameter is
$\Sigma^{\text{KL}}_{ii}=\Lambda_{ii}^{-1}\quad\text{for}\quad i=1,\dots,d.$
Thus, the precision is matched by the variational density, but the variance is underestimated. Lemma 1 presents the weighted Fisher divergence for a general weight matrix $M$, which is $I_{d}$ in the Fisher divergence and $\Sigma$ in the score-based divergence.
Lemma 1.
The $M$-weighted Fisher divergence between a Gaussian target $p(\theta|y)={\text{N}}(\theta|\nu,\Lambda^{-1})$ and Gaussian variational approximation $q(\theta)={\text{N}}(\theta|\mu,\Sigma)$ is
$S_{M}(q\|p)$ $=\text{\rm tr}(\Sigma^{-1}M)+\text{\rm tr}(\Lambda M\Lambda\Sigma)-2\text{\rm tr}(M\Lambda)$
$\quad+(\mu-\nu)^{\top}\Lambda M\Lambda(\mu-\nu).$
If $\Sigma$ is a diagonal matrix, then
$S_{M}(q\|p)$ $=\sum_{i=1}^{d}\{\Sigma_{ii}^{-1}M_{ii}+(\Lambda M\Lambda)_{ii}\Sigma_{ii}\}-2\text{\rm tr}(M\Lambda)$
$\quad+(\mu-\nu)^{\top}\Lambda M\Lambda(\mu-\nu).$
Proof.
The $M$-weighted Fisher divergence is
$S_{M}(q\|p)$ $={\text{E}}_{q}\{\|\nabla_{\theta}\log q(\theta)-\nabla_{\theta}\log p(\theta|y)\|_{M}^{2}\}$
$={\text{E}}_{q}\{\|\Sigma^{-1}(\theta-\mu)-\Lambda(\theta-\nu)\|_{M}^{2}\}$
$={\text{E}}_{q}\{\|(\Sigma^{-1}-\Lambda)(\theta-\mu)-\Lambda(\mu-\nu)\|_{M}^{2}\}$
$=\text{\rm tr}\{(\Sigma^{-1}-\Lambda)M(\Sigma^{-1}-\Lambda)\Sigma\}$
$\quad+(\mu-\nu)^{\top}\Lambda M\Lambda(\mu-\nu)$
$=\text{\rm tr}(\Sigma^{-1}M)+\text{\rm tr}(\Lambda M\Lambda\Sigma)-2\text{\rm tr}(M\Lambda)$
$\quad+(\mu-\nu)^{\top}\Lambda M\Lambda(\mu-\nu).$
The final result arises from $\text{\rm tr}(AB)=\sum_{i=1}^{d}A_{ii}B_{ii}$ if $A$ is a diagonal matrix. ∎
From Lemma 1, $\nabla_{\mu}S_{M}(q\|p)=2\Lambda M\Lambda(\mu-\nu)$. Thus, $\nabla_{\mu}S_{M}(q\|p)=0$ implies $\mu=\nu$, and the true posterior mean is recovered for any $M$-weighted Fisher divergence where $M$ is independent of $\mu$. Under the mean-field assumption, at this optimal value of $\mu$,
$S_{M}(q\|p)=\sum_{i=1}^{d}\{\Sigma_{ii}^{-1}M_{ii}+(\Lambda M\Lambda)_{ii}\Sigma_{ii}\}-2\text{\rm tr}(M\Lambda).$
If the weight $M$ is independent of $\Sigma$, then $\nabla_{\Sigma_{ii}}S_{M}(q\|p)=(\Lambda M\Lambda)_{ii}-M_{ii}/\Sigma_{ii}^{2}=0$ implies
$\Sigma_{ii}=\sqrt{M_{ii}/(\Lambda M\Lambda)_{ii}}\quad\text{for}\quad i=1,\dots,d.$
Thus a closed form solution exists for any $M$ independent of $\Sigma$. Moreover, if $M$ is a diagonal matrix, then
$\Sigma_{ii}=\sqrt{\frac{M_{ii}}{\sum_{j=1}^{d}M_{jj}\Lambda_{ij}^{2}}}\quad\text{for}\quad i=1,\dots,d.$ (1)
When $M_{ii}=1$ $\forall$ $i$, we recover the Fisher divergence for which the optimal variance parameters are
$\Sigma^{\text{F}}_{ii}=\frac{1}{\sqrt{\sum_{j=1}^{d}\Lambda_{ij}^{2}}}\quad\text{for}\quad i=1,\dots,d.$ (2)
Optimal variational parameters for score-based divergence under mean-field assumption have been presented in Margossian et al. (2024), and a discussion is included here for completeness. Plugging $M=\Sigma$ in Lemma 1,
$S(q\|p)$ $=d+\sum_{i=1}^{d}\sum_{j=1}^{d}\Sigma_{ii}\Sigma_{jj}\Lambda_{ij}^{2}-2\sum_{i=1}^{d}\Sigma_{ii}\Lambda_{ii},$
at the optimal value of $\mu$. Let $\mathbf{s}=(s_{1},\dots,s_{d})^{\top}$ such that $s_{i}=\Sigma_{ii}\Lambda_{ii}\geq 0$, and $H$ be a $d\times d$ symmetric matrix with $H_{ij}=\Lambda_{ij}^{2}/(\Lambda_{ii}\Lambda_{jj})$. Then $S(q\|p)=d+2F(\mathbf{s})$, where
$F(\mathbf{s})=\frac{1}{2}\mathbf{s}^{\top}H\mathbf{s}-\mathbf{1}^{\top}\mathbf{s}.$ (3)
Thus the optimal $\Sigma^{\text{S}}_{ii}$ that minimizes $S(q\|p)$ can be obtained by solving a non-negative quadratic program (NQP) for $\mathbf{s}$. NQP is the problem of minimizing the quadratic objective function in (3) subject to the constraint $s_{i}\geq 0$ $\forall$ $i$. Since $\Lambda$ is positive definite,
$x^{\top}Hx=\sum_{i=1}^{d}\sum_{j=1}^{d}(x_{i}/\Lambda_{ii})\Lambda_{ij}^{2}(x_{j}/\Lambda_{jj})=y^{\top}\Lambda y>0$
for any $x=(x_{1},\dots,x_{d})^{\top}\in\mathbb{R}^{d}$ and $y=(x_{1}/\Lambda_{11},\dots,x_{d}/\Lambda_{dd})^{\top}$. Thus $H$ is symmetric positive definite, which implies that $F(\mathbf{s})$ is bounded below and its optimization is convex. However, there is no analytic solution for the global minimum due to the non-negativity constraints and iterative solutions are required (Sha et al., 2003). The Karush-Kuhn-Tucker (KKT) conditions are first derivative tests that can be used to check whether a solution returned by an iterative solver is indeed a local optimum. For the NQP in (3), the KKT conditions state that $\forall$ $i=1,\dots,d$, either (a) $s_{i}=0$ and $(Hs)_{i}>1$ or (b) $s_{i}>0$ and $(Hs)_{i}=1$. Note that $\nabla_{\mathbf{s}}F(\mathbf{s})=H\mathbf{s}-{\bf 1}$. These conditions correspond to cases where the constraint is active or inactive at the optimum. Case (a) implies $\Sigma_{ii}^{\text{S}}=0$, meaning that the variational density collapses to a point estimate in the $i$th dimension. Note that the KL and Fisher divergences do not face this issue of “variational collapse”. Case (b) implies
$(H\mathbf{s})_{i}=$ $\sum_{j=1}^{d}H_{ij}s_{j}=\sum_{j=1}^{d}\frac{\Lambda_{ij}^{2}}{\Lambda_{ii}\Lambda_{jj}}\Sigma_{jj}\Lambda_{jj}=1$ (4)
$\implies\sum_{j=1}^{d}\Lambda_{ij}^{2}\Sigma_{jj}=\Lambda_{ii}.$
Figure 1: Variance parameter comparisons for Fisher and score-based divergence.
Next, we investigate how the variance parameters $\{\Sigma_{ii}\}$ obtained by minimizing the weighted Fisher divergence compare to those obtained by minimizing the KLD.
Theorem 1.
Suppose the target is a multivariate Gaussian with non-diagonal precision matrix $\Lambda$, and the variational family is Gaussian with diagonal covariance matrix $\Sigma$. Let $\Sigma^{\text{KL}}_{ii}$, $\Sigma^{\text{M}}_{ii}$ and $\Sigma^{\text{S}}_{ii}$ denote the optimal value of the $i$th diagonal element of $\Sigma$ obtained by minimizing the KL, $M$-weighted Fisher and score-based divergences respectively, where $M$ is a positive definite diagonal matrix independent of $\Sigma$. Then
$\Sigma^{\text{M}}_{ii}\leq\Sigma^{\text{KL}}_{ii}\quad\text{and}\quad\Sigma^{\text{S}}_{ii}\leq\Sigma^{\text{KL}}_{ii}\quad\text{for}\quad i=1,\dots,d.$
Moreover, $\exists\,i\in\{1,\dots,d\}$ such that $\Sigma_{ii}^{M}<\Sigma_{ii}^{KL}$ and $\Sigma_{ii}^{S}<\Sigma_{ii}^{KL}$.
Proof.
We first prove $\Sigma^{\text{M}}_{ii}\leq\Sigma^{\text{KL}}_{ii}$ $\forall\,i$. From (1),
$\Sigma^{\text{M}}_{ii}=\sqrt{\frac{M_{ii}}{\sum_{j=1}^{d}M_{jj}\Lambda_{ij}^{2}}}\leq\sqrt{\frac{M_{ii}}{M_{ii}\Lambda_{ii}^{2}}}=\frac{1}{\Lambda_{ii}}=\Sigma_{ii}^{\text{\rm KL}}.$ (5)
Since $\Lambda$ has at least one nonzero off-diagonal entry, $\exists\,i\in\{1,\dots,d\}$ such that the inequality in (5) is strict. The proof for $\Sigma^{\text{S}}_{ii}\leq\Sigma^{\text{KL}}_{ii}$ is given in Margossian et al. (2024) and we include it here for entirety. From the KKT conditions discussed earlier, if case (a) applies, then $\Sigma^{\text{S}}_{ii}=0<\Sigma^{\text{KL}}_{ii}$. Otherwise, case (b) applies and (4) implies that
$\Lambda_{ii}^{2}\Sigma^{\text{S}}_{ii}\leq\sum_{j}\Lambda_{ij}^{2}\Sigma^{\text{S}}_{ii}=\Lambda_{ii}\implies\Sigma^{\text{S}}_{ii}\leq\frac{1}{\Lambda_{ii}}=\Sigma^{\text{KL}}_{ii}.$ (6)
To obtain the strict inequality, note that if case (a) applies for at least one $i$, then $\Sigma_{ii}^{S}<\Sigma_{ii}^{KL}$ for such an $i$. Otherwise, case (b) applies $\forall$ $i$. Since $\Lambda$ has at least one nonzero off-diagonal entry, $\exists\,i\in\{1,\dots,d\}$ such that the first inequality in (6) is strict. ∎
From Theorem 1, both the weighted Fisher and score-based divergences tend to underestimate the posterior variance more severely than KLD under the mean-field assumption, but ordering between the Fisher and score-based divergences is more nuanced. If case (a) of the KKT conditions apply, then $\Sigma^{\text{S}}_{ii}=0<\Sigma^{\text{F}}_{ii}$. If case (b) applies, then from (2) and (4),
$\Lambda_{ii}^{2}\Sigma^{\text{S}}_{ii}$ $\leq\sum_{j=1}^{d}\Lambda_{ij}^{2}\Sigma^{\text{S}}_{jj}=\Lambda_{ii}\Sigma^{\text{F}}_{ii}\sqrt{\sum\nolimits_{j=1}^{d}\Lambda_{ij}^{2}}$
$\implies\Sigma^{\text{S}}_{ii}\leq\Sigma^{\text{F}}_{ii}\frac{\sqrt{\sum_{j=1}^{d}\Lambda_{ij}^{2}}}{\Lambda_{ii}}.$
Moreover, if $\Lambda$ is a diagonally dominant matrix such that $\sum_{j\neq i}|\Lambda_{ij}|\leq|\Lambda_{ii}|\;\forall\;i$, then
$\Sigma^{\text{S}}_{ii}$ $\leq\Sigma^{\text{F}}_{ii}\frac{\sqrt{\Lambda_{ii}^{2}+\sum_{j\neq i}\Lambda_{ij}^{2}}}{\Lambda_{ii}}\leq\Sigma^{\text{F}}_{ii}\frac{\sqrt{\Lambda_{ii}^{2}+(\sum_{j\neq i}|\Lambda_{ij}|)^{2}}}{\Lambda_{ii}}$
$\leq\Sigma^{\text{F}}_{ii}\frac{\sqrt{\Lambda_{ii}^{2}+\Lambda_{ii}^{2}}}{\Lambda_{ii}}=\sqrt{2}\Sigma^{\text{F}}_{ii}.$
Thus the ratio of $\Sigma^{\text{S}}_{ii}/\Sigma^{\text{F}}_{ii}$ is bounded by $\sqrt{2}$ $\forall\;i$ if $\Lambda$ is diagonally dominant.
For a more concrete comparison of posterior variance approximation based on Fisher and score-based divergences, consider a three-dimensional Gaussian target with precision matrix,
$\Lambda=\begin{bmatrix}1&a&b\\ a&1&c\\ b&c&1\end{bmatrix}.$
For Fisher divergence, $\Sigma_{ii}^{\text{F}}$ can be obtained from (2), while the splitting conic solver (SCS, O’Donoghue et al., 2016) in the CVXPY Python package is used to solve the NQP in (3) for score-based divergence. SCS is designed for convex optimization problems characterized by conic constraints, such as nonnegative constraints. It decomposes the optimization task into smaller subproblems solved iteratively by operator-splitting techniques.
Fig 1 illustrates how variance parameters obtained from the Fisher and score-based divergences compare by varying the conditional correlations $a$, $b$ and $c$, with each plot representing a value of $c$. The colored regions represent configurations for which $\Lambda$ is positive definite. The red, blue and green regions indicate where $\Sigma_{ii}^{S}\leq\Sigma_{ii}^{F}$ for all cases, only two cases and only one case respectively. There is no region such that $\Sigma_{ii}^{S}>\Sigma_{ii}^{F}$ $\forall$ $i$. Variance estimates based on score-based divergence have a higher tendency to exceed those based on Fisher divergence when the conditional correlation $a$, $b$ or $c$ has a large magnitude. The orange-bordered region indicates where $\Lambda$ is diagonally dominant, and $\Sigma^{\text{S}}_{ii}/\Sigma^{\text{F}}_{ii}$ can be bounded even more tightly by 1 instead of only $\sqrt{2}$ in this example.
3 Ordering of divergences for non-Gaussian target
Figure 2: Gaussian variational approximations for Student’s $t$ and log transformed inverse gamma.
Next, we study ordering of the Fisher, score-based and KL divergences in posterior mode and variance estimation when the target $p(\theta|y)$ is a univariate non-Gaussian density, while the variational density $q(\theta)$ is ${\text{N}}(\mu,\sigma^{2})$ with mean $\mu\in\mathbb{R}$ and standard deviation $\sigma>0$. In the examples below, the VI objective function is available in closed form or can be computed numerically, and optimal variational approximations are obtained using L-BFGS via the optim function in R. For univariate densities, $S(q\|p)=\sigma^{2}F(q\|p)$, so we only present the Fisher divergence. All derivations are given in the supplement.
Let $m_{*}$ and $\sigma^{2}_{*}$ denote the mode and variance of the target density. To evaluate the performance of different divergences, we use three criteria, normalized absolute difference in mode: $|\mu-m_{*}|/\sigma_{*}$, variance ratio: $\sigma^{2}/\sigma^{2}_{*}$, and integrated absolute error : $\text{IAE}(q)=\int|q(\theta)-p(\theta|y)|d\theta$, which is invariant under monotone transformations of $\theta$ and lies in [0, 2]. We define $\text{Accuracy}(q)=1-\text{IAE}(q)/2$, where a higher accuracy indicates a better approximation of the target.
3.1 Student’s $t$
First, we consider the univariate Student’s $t$ as target, which is symmetric but has heavier tails than a Gaussian. The probability density function (pdf) for $\theta\sim t(\nu)$ is $p(\theta|y)=\left(1+\theta^{2}/\nu\right)^{-(\nu+1)/2}\Gamma(\frac{\nu+1}{2})/(\sqrt{\pi\nu}\,\Gamma(\frac{\nu}{2}))$, where $\nu\in\{3,5,10\}$ is the degrees of freedom. For KLD, we maximize the lower bound,
${\mathcal{L}}$ $=\log\left\{\Gamma\left(\frac{\nu+1}{2}\right)\right\}-\frac{1}{2}\log\left\{\Gamma^{2}\left(\frac{\nu}{2}\right)\nu\right\}$
$\quad-\frac{\nu+1}{2}{\text{E}}_{q}\log\left(1+\frac{\theta^{2}}{\nu}\right)+\frac{1}{2}\log(2\sigma^{2})+\frac{1}{2}.$
The Fisher divergence is
$F(q\|p)$ $=(\nu+1)^{2}{\text{E}}_{q}\left\{\frac{\theta^{2}}{(\nu+\theta^{2})^{2}}\right\}+\frac{1}{\sigma^{2}}$
$\quad-\frac{2(\nu+1)}{\sigma^{2}}{\text{E}}_{q}\left\{\frac{\theta(\theta-\mu)}{\nu+\theta^{2}}\right\}.$
All three divergences successfully capture the mode of the target at 0. From Table 1, SD exhibits the most posterior variance underestimation, followed by FD, while KLD has the least underestimation. In terms of the IAE, both FD and SD yield approximations with higher accuracy than KLD. Fig 2 (first 3 plots) compares variational densities with the target. KLD tends to underestimate the probability mass around the mode more severely than FD and SD.
$\sigma^{2}/\sigma^{2}_{*}$ accuracy
$\nu$ KLD FD SD KLD FD SD
3 0.529 0.428 0.372 92.18 93.66 92.62
5 0.818 0.728 0.681 94.72 95.82 95.97
10 0.950 0.909 0.889 97.01 97.55 97.73
Table 1: Variance ratio and accuracy for Student’s $t$.
Figure 3: Gaussian variational approximations for skew normal.
$|\mu-m_{*}|/\sigma_{*}$ $\sigma^{2}/\sigma^{2}_{*}$ accuracy
$(t,\lambda)$ KLD FD SD KLD FD SD KLD FD SD
(1, 1) 0.07 0.07 0.07 0.992 0.984 0.979 98.27 98.31 98.32
(1, 2) 0.25 0.23 0.20 0.919 0.851 0.803 93.77 93.81 93.79
(1, 5) 0.66 0.91 0.07 0.677 0.642 0.248 83.93 76.44 68.50
(5, 1) 0.66 0.91 0.07 0.677 0.642 0.248 83.92 76.42 68.49
(5, 2) 0.94 2.20 0.10 0.504 0.757 0.054 76.50 45.38 37.06
(5, 5) 1.20 2.94 0.09 0.352 0.644 0.008 68.00 30.35 16.35
Table 2: Results of KL, Fisher and score-based divergences for skew normal.
3.2 Log transformed inverse-gamma
Next, consider the normal sample model in Tan and Chen (2024), where $y_{i}|\theta\sim{\text{N}}(0,\exp(\theta))$ for $i=1,\dots,n$, with prior $\exp(\theta)\sim{\text{IG}}(a_{0},b_{0})$ and $a_{0}=b_{0}=0.01$. The true posterior of $\exp(\theta)$ is ${\text{IG}}(a_{1},b_{1})$, where $a_{1}=a_{0}+n/2$ and $b_{1}=b_{0}+\sum_{i=1}^{n}y_{i}^{2}/2$. We simulate $n=6$ observations by setting $\exp(\theta)=225$. The lower bound for minimizing the KLD is
${\mathcal{L}}$ $=\frac{1-n}{2}\log(2\pi)+a_{0}\log b_{0}-\log\Gamma(a_{0})-a_{1}\mu$
$\quad-\exp\left(\frac{\sigma^{2}-2\mu}{2}\right)T+\frac{1}{2}\log(\sigma^{2})+\frac{1}{2}.$
The Fisher divergence is
$F(q\|p)$ $=a_{1}^{2}+b_{1}^{2}\exp\left(2\sigma^{2}-2\mu\right)$
$\quad-2b_{1}\left(a_{1}+1\right)\exp(\sigma^{2}/2-\mu)+{1}/{\sigma^{2}}.$
Table 3 shows that SD provides the most accurate mode approximation but also exhibits the most severe variance underestimation. On the other hand, KLD yields the least accurate mode approximation while most closely matching the target variance. Overall, KLD has the highest accuracy followed by FD and SD. A visualization is given in Fig 2 (last plot).
KLD FD SD
$|\mu-m_{*}|/\sigma_{*}$ 0.26 0.23 0.18
$\sigma^{2}/\sigma^{2}_{*}$ 0.845 0.732 0.674
accuracy 92.67 91.91 91.53
Table 3: Results for log transformed inverse-gamma.
3.3 Skew normal
Finally, let the target be a univariate skew normal. The pdf of $\theta\sim{\text{SN}}(m,t,\lambda)$ is $p(\theta|y)=2\,\phi(\theta|m,t^{2})\,\Phi\{\lambda(\theta-m)\}$, where $m\in\mathbb{R}$, $t>0$ and $\lambda\in\mathbb{R}$ are the location, scale and skewness parameters respectively, and $\Phi(\cdot)$ is cumulative distribution function (cdf) of the standard normal. For KLD, we maximize the lower bound,
${\mathcal{L}}$ $=\log 2-\log(t)-\frac{\sigma^{2}+(\mu-m)^{2}}{2t^{2}}$
$\quad+{\text{E}}_{q}\log[\Phi\{\lambda(\theta-m)\}]+\log(\sigma)+\frac{1}{2}.$
The Fisher divergence is
$F(q\|p)$ $=\frac{\sigma^{2}+(\mu-m)^{2}}{t^{4}}+\lambda^{2}{\text{E}}_{q}\left[\frac{\phi^{2}\{\lambda(\theta-m)\}}{\Phi^{2}\{\lambda(\theta-m)\}}\right]$
$\quad-\frac{2\lambda}{t^{2}}\,{\text{E}}_{q}\left[(\theta-m)\frac{\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}\right]-\frac{2}{t^{2}}$
$\quad+\frac{2\lambda}{\sigma^{2}}{\text{E}}_{q}\left[(\theta-\mu)\frac{\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}\right]+\frac{1}{\sigma^{2}}.$
We set $m=0$ and let $t\in\{1,5\}$ and $\lambda\in\{1,2,5\}$. From Table 2, SD captured the mode most accurately across all settings, followed by KLD and FD (for which estimation of the mode is very poor when both scale and skewness are large). All three divergences underestimated the variance, with KLD and FD exhibiting less underestimation than SD (for which the variance estimate approaches zero as $t$ and $\lambda$ increase). The accuracies of FD and SD are very close to KLD when $t=1$ and $\lambda\in\{1,2\}$. However, KLD has higher accuracy than both FD and SD as skewness and scale increase. Fig 3 shows density plots of the target and optimal variational approximations. SD is good at identifying the mode, whereas FD and KLD estimate the variance more accurately. Multiple local minimums were detected for the SD in this context.
The examples in this section suggest that SD captures the mode more accurately than FD and KLD if the target density is skewed, but underestimates the posterior variance most severely. KLD has higher accuracy than both FD and SD when the target is skewed, but lower accuracy when the target is symmetric and has heavy tails.
4 Sparse Gaussian variational approximations
Next, we consider Gaussian VI for hierarchical models and compare performances of the Fisher, score-based and KL divergences. Given observed data $y=(y_{1},\dots,y_{n})^{\top}$, the variable $\theta=(\theta_{L}^{\top},\theta_{G}^{\top})^{\top}\in\mathbb{R}^{d}$ of a two-tier hierarchical model can be partitioned into a global variable $\theta_{G}$ that is shared among all observations and local variables $\theta_{L}=(b_{1}^{\top},\dots,b_{n}^{\top})^{\top}$, where $b_{i}$ is specific to the observation $y_{i}$ for $i=1,\dots,n$. Let the joint likelihood of the model be
$p(y,\theta)=p(\theta_{G})p(b_{1},\dots,b_{\ell}|\theta_{G})\times\\ \prod_{k=\ell+1}^{n}p(b_{k}|b_{k-1},\dots,b_{k-\ell},\theta_{G})\prod_{i=1}^{n}p(y_{i}|\theta_{G},b_{i}),$ (7)
where $\{y_{i}\}$ are conditionally independent given $\theta$, and the $\{b_{i}\}$ follow an $\ell$th order Markov model given $\theta_{G}$. Thus $\{b_{i}\}$ are conditionally independent of each other a posteriori given the $\ell$ neighboring values and $\theta_{G}$. In a random effects model, $\{b_{i}\}$ are the random effects with $\ell=0$, while for a state space model, $\{b_{i}\}$ are the latent states with $\ell=1$.
Let $q_{\lambda}(\theta)$ be ${\text{N}}(\mu,\Sigma)$, a Gaussian variational approximation of the posterior density with mean $\mu\in\mathbb{R}^{d}$ and covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$. Consider a Cholesky decomposition of the precision matrix $\Omega=\Sigma^{-1}=TT^{\top}$ where $T$ is a lower triangular matrix, and denote the variational parameters as $\lambda=(\mu^{\top},\text{vech}(T)^{\top})^{\top}$, where $\text{vech}(\cdot)$ is an operator that stacks lower triangular elements of a matrix columnwise from left to right into a vector.
In a multivariate Gaussian, conditional independence implies sparse structure in the precision matrix, with $\Omega_{ij}=0$ if $\theta_{i}$ and $\theta_{j}$ are conditionally independent given the remaining variables. By Proposition 1 of Rothman et al. (2010), Cholesky factor $T$ has the same row-banded structure as $\Omega$. Suppose $T$ is block partitioned according to $(b_{1}^{\top},\dots,b_{n}^{\top},\theta_{G}^{\top})^{\top}$, with corresponding blocks $T_{ij}$ for $i,j=1,\dots,n+1$. If $T_{ij}=0$ for $1\leq j\leq i-l$ and $j>i$ (as $T$ is lower-triangular), then $\Omega$ reflects the conditional independence structure of the true posterior. When $\theta$ is high-dimensional, exploiting the conditional independence structure in the model is essential to making Gaussian VI feasible, as the number of parameters to be optimized in $T$ grows quadratically with $n$. However, after imposing sparsity on $T$, the number of parameters only grows linearly with $n$.
We propose two SGD approaches for minimizing the Fisher and score-based divergences, as SGD allows updating of only the elements in $T$ that are not constrained to zero by sparsity. Algorithms that minimize the FD and SD based on unbiased gradient estimates derived using the reparametrization trick are named FDr and SDr (“r” for reparametrization trick, see Table 4), while algorithms that rely on a batch approximation of the FD and SD are named FDb and SDb (“b” for batch approximation, see Table 5) respectively.
5 SGD based on reparameterization trick
The Fisher and score-based divergences between $q_{\lambda}(\theta)$ and $p(\theta|y)$ can be written as
$F(\lambda)$ $={\text{E}}_{q}[g(\lambda,\theta)^{\top}g(\lambda,\theta)],$
$S(\lambda)$ $={\text{E}}_{q}[f(\lambda,\theta)^{\top}f(\lambda,\theta)],$
respectively, where $g(\lambda,\theta)=\nabla_{\theta}\log h(\theta)-\nabla_{\theta}\log q_{\lambda}(\theta)$, $f(\lambda,\theta)=T^{-1}g(\lambda,\theta)$ and $h(\theta)=p(\theta)p(y|\theta)$ as defined previously. We derive gradients of the Fisher and score-based divergences by applying the reparametrization trick of Kingma and Welling (2014). Instead of simulating $\theta$ directly from $q_{\lambda}(\theta)$, we generate $z\sim{\text{N}}(0,I_{d})$ and compute $\theta=\mu+T^{-\top}z$. Thus
$F(\lambda)$ $={\text{E}}_{\phi}\left\{g(\lambda,\mu+T^{-\top}z)^{\top}g(\lambda,\mu+T^{-\top}z)\right\},$
$S(\lambda)$ $={\text{E}}_{\phi}\left\{f(\lambda,\mu+T^{-\top}z)^{\top}f(\lambda,\mu+T^{-\top}z)\right\},$
where ${\text{E}}_{\phi}(\cdot)$ denotes expectation with respect to $\phi(z)$, the pdf of ${\text{N}}(0,I_{d})$. Note that
$g(\lambda,\theta)$ $=\nabla_{\theta}\log h(\theta)+TT^{\top}(\theta-\mu),$
$f(\lambda,\theta)$ $=T^{-1}\nabla_{\theta}\log h(\theta)+T^{\top}(\theta-\mu),$
both of which depends on $\lambda$ directly as well as through $\theta$. Applying the chain rule,
$\nabla_{\mu}F(\lambda)$ $=2{\text{E}}_{\phi}\{\nabla_{\theta}^{2}\log h(\theta)g(\lambda,\theta)\},$
$\nabla_{\mu}S(\lambda)$ $=2{\text{E}}_{\phi}\{\nabla_{\theta}^{2}\log h(\theta)\Sigma g(\lambda,\theta)\},$
$\nabla_{\text{\rm vech}(T)}F(\lambda)$ $=2{\text{E}}_{\phi}\text{\rm vech}\left\{g(\lambda,\theta)z^{\top}\right.$
$\left.\quad-T^{-\top}zg(\lambda,\theta)^{\top}\nabla_{\theta}^{2}\log h(\theta)T^{-\top}\right\},$
$\nabla_{\text{\rm vech}(T)}S(\lambda)$ $=-2{\text{E}}_{\phi}\text{\rm vech}\left\{\Sigma g(\lambda,\theta)\nabla_{\theta}\log h(\theta)^{\top}T^{-\top}\right.$
$\left.\quad+T^{-\top}zg(\lambda,\theta)^{\top}\Sigma\nabla_{\theta}^{2}\log h(\theta)T^{-\top}\right\}.$
Detailed derivations are given in the supplement.
Unbiased estimates of the gradients can be obtained by sampling from $\phi(z)$. All gradient computations can be done efficiently even in high-dimensions, as they only involve sparse matrix multiplications and solutions of sparse triangular linear systems. The Hessian $\nabla_{\theta}^{2}\log h(\theta)$ has the same block sparse structure as $\Omega$, as $b_{i}$ and $b_{j}$ only occur in the same factor of (7) if $b_{j}$ is one of the $\ell$ neighboring values of $b_{i}$. For $\nabla_{\text{\rm vech}(T)}F(\lambda)$ and $\nabla_{\text{\rm vech}(T)}S(\lambda)$, we only need to compute elements corresponding to those in $\text{\rm vech}(T)$ that are not fixed by sparsity. For instance, to compute the second term in $\nabla_{\text{\rm vech}(T)}F(\lambda)$, we just find $u=T^{-\top}z$ and $v=T^{-1}\nabla_{\theta}^{2}h(\theta)g(\lambda,\theta)$, and then form $u_{i}v_{j}$ for nonzero elements $(i,j)$ of $T$.
The update for $T$ in SGD does not ensure positiveness of the diagonal entries. Hence, we introduce $T^{*}$ such that $T^{*}_{ii}=\log(T_{ii})$ for $i=1,\dots,n,$ and $T^{*}_{ij}=T_{ij}\;for\;i\neq j$. Let $J$ be a $d\times d$ matrix with diagonal equal to $\text{\rm diag}(T)$ and all off-diagonal entries being 1, and $D$ be a diagonal matrix with the diagonal given by $\text{\rm vech}(J)$. Then $\nabla_{\text{\rm vech}(T^{*})}F(\lambda)=D\nabla_{\text{\rm vech}(T)}F(\lambda)$ and updates for $T^{*}$ are unconstrained.
KLD FDr and SDr Initialize $\mu$ and $T$. For $t=1,2,\dots N$, 1. Generate $z\sim{\text{N}}(0,I_{d})$. 2. Compute $u=T^{-\top}z$, $\theta=\mu+u$ and $g=\nabla_{\theta}\log h(\theta)+Tz$. 3. $\mu\leftarrow\mu+\rho_{t}g$. 4. $v=T^{-1}g$. 5. $g_{T}=-uv^{\top}$. 6. $T^{*}\leftarrow T^{*}+\rho_{t}\;Dg_{T}$. 7. Obtain $T$ from $T^{*}$. 3. If SD, $g\leftarrow T^{-1}g$. $z\leftarrow z-g$ $g\leftarrow T^{-\top}g$. 4. $w=\nabla_{\theta}^{2}\log h(\theta)g$,
$v=T^{-1}w$.
5. $\mu\leftarrow\mu-2\rho_{t}w$. 6. $g_{T}=gz^{\top}-uv^{\top}$. 7. $T^{*}\leftarrow T^{*}-2\rho_{t}Dg_{T}$. 8. Obtain $T$ from $T^{*}$.
Table 4: SGD algorithms based on reparametrization trick.
Table 4 outlines the SGD algorithms for updating $(\mu,T)$ by minimizing the Fisher, score-based or KL (derived in Tan and Nott, 2018) divergences. The stepsize $\rho_{t}$ is computed elementwise adaptively using Adadelta (Zeiler, 2012). All three algorithms compute $g(\lambda,\theta)$, but the KLD based algorithm uses $g(\lambda,\theta)$ directly to update $\mu$ and $T$, while the FD and SD based algorithms use the Hessian $\nabla_{\theta}^{2}\log h(\theta)$ to condition $g(\lambda,\theta)$ and are hence more computationally intensive.
5.1 Analysis of variance of gradient estimates
We study variance of the unbiased gradient estimates derived by applying the reparametrization trick on the KL, Fisher and score-based divergences, to examine effects of the conditioning matrix, $\nabla_{\theta}^{2}\log h(\theta)$. The variance of these gradients plays a crucial role in the stability of the optimization, as large variance can cause a zigzag phenomenon, making convergence difficult. For a closed form analysis, we assume the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$. Then $\nabla_{\theta}\log h(\theta)=-\Lambda(\theta-\nu)$ and $\nabla_{\theta}^{2}\log h(\theta)=-\Lambda$.
Figure 4: First two plots show the true precision matrix $\Lambda$ and mean $\nu$, and the third plot contains boxplots of the standard deviation (s.d.) in gradient estimates for $\{\mu_{i}\}$ and $\{T_{ii}\}$.
From Table 4, gradient estimates with respect to $\mu$ for the KL, Fisher and score-based divergences based on a single sample are
$g_{\mu}^{\text{\rm KL}}=Az-\Lambda(\mu-\nu),\quad g_{\mu}^{\text{F}}=2\Lambda g_{\mu}^{\text{\rm KL}},\quad g_{\mu}^{\text{S}}=2\Lambda\Sigma g_{\mu}^{\text{\rm KL}},$
where $A=T-\Lambda T^{-\top}$. The stochasticity stems from drawing $z\sim{\text{N}}(0,I_{d})$ and $\text{\rm Var}(g_{\mu}^{\text{\rm KL}})=AA^{\top}$, while
$\text{\rm Var}(g_{\mu}^{\text{F}})=4\Lambda\text{\rm Var}(g_{\mu}^{\text{\rm KL}})\Lambda,\;\;\text{\rm Var}(g_{\mu}^{\text{S}})=4\Lambda\Sigma\text{\rm Var}(g_{\mu}^{\text{\rm KL}})\Sigma\Lambda.$
Similarly, from Table 4, the gradient estimates with respect to $T$ are
$g_{T}^{\text{\rm KL}}$ $=T^{-\top}z(\mu-\nu)^{\top}\Lambda T^{-\top}-T^{-\top}zz^{\top}A^{\top}T^{-\top},$
$g_{T}^{\text{F}}$ $=2\{\Lambda(\mu-\nu)z^{\top}+T^{-\top}z(\mu-\nu)^{\top}\Lambda^{2}T^{-\top}$
$\quad-Azz^{\top}-T^{-\top}zz^{\top}A^{\top}\Lambda T^{-\top}\},$
$g_{T}^{\text{S}}$ $=2[\Sigma\Lambda(\mu-\nu)\{z^{\top}T^{-1}+(\mu-\nu)^{\top}\}\Lambda T^{-\top}$
$\quad-\Sigma A\{zz^{\top}T^{-1}+z(\mu-\nu)^{\top}\}\Lambda T^{-\top}$
$\quad+T^{-\top}\{z(\mu-\nu)^{\top}\Lambda-zz^{\top}A^{\top}\}\Sigma\Lambda T^{-\top}].$
The variance of these gradient estimates depends on the mean and precision of the true target, which is fixed, and that of the variational approximation, which keeps changing during SGD. Suppose $\Lambda$ and $T$ are both diagonal matrices, then
$\text{\rm Var}(g_{\mu_{i}}^{\text{\rm KL}})$ $=T_{ii}^{2}-2\Lambda_{ii}+\Lambda_{ii}^{2}T_{ii}^{-2},$
$\text{\rm Var}(g_{\mu_{i}}^{\text{F}})$ $=4\Lambda_{ii}^{2}\text{\rm Var}(g_{\mu_{i}}^{\text{\rm KL}}),$
$\text{\rm Var}(g_{\mu_{i}}^{\text{S}})$ $=(4\Lambda_{ii}^{2}/T_{ii}^{4})\,\text{\rm Var}(g_{\mu_{i}}^{\text{\rm KL}}),$
$\text{\rm Var}(g_{T_{ii}}^{\text{\rm KL}})$ $=T_{ii}^{-4}\left\{\Lambda_{ii}^{2}(\mu_{i}-\nu_{i})^{2}+2\left(T_{ii}-\Lambda_{ii}/T_{ii}\right)^{2}\right\},$
$\text{\rm Var}(g_{T_{ii}}^{\text{F}})$ $=4(T_{ii}^{2}+\Lambda_{ii})^{2}\text{\rm Var}(g_{T_{ii}}^{\text{\rm KL}}).$
$\text{\rm Var}(g_{T_{ii}}^{\text{S}})$ $=4\Lambda_{ii}^{2}T_{ii}^{-8}\left\{\left(3\Lambda_{ii}-T_{ii}^{2}\right)^{2}(\mu_{i}-\nu_{i})^{2}\right.$
$\quad+\left.8\left(T_{ii}-\Lambda_{ii}/T_{ii}\right)^{2}\right\}.$
It can be verified that these variances are zero at full convergence, when $\mu_{i}=\nu_{i}$ and $T_{ii}^{2}=\Lambda_{ii}$ $\forall$ $i$. The variance of gradients with respect to $\mu$ of FD and SD are larger than that of KLD if $\Lambda_{ii}>0.5$ and $\Lambda_{ii}/T_{ii}^{2}>0.5$ respectively. Assuming $\mu_{i}=\nu_{i}$ for the SD, the variance of gradients with respect to $T$ of FD and SD are larger than that of KLD if $T_{ii}^{2}+\Lambda_{ii}>0.5$ and $\Lambda_{ii}/T_{ii}^{2}>0.25$ respectively.
In summary, the variance of gradient estimates based on FD is larger than that of KL once $\Lambda_{ii}>0.5$ regardless of the values of the variational parameters, and variance inflation is larger for $T$ than $\mu$. For SD, the inflation factor involves the ratio $\Lambda_{ii}/T_{ii}^{2}$, so variance inflation relative to KLD can be reduced if $T_{ii}^{2}>\Lambda_{ii}$.
Next, we investigate the variance of gradient estimates for a multivariate Gaussian target with $d=49$ in a real setting. The true precision matrix $\Lambda$ and mean $\nu$, visualized in the first two plots of Fig 4, are derived from MCMC samples obtained by fitting a logistic regression model to the German credit data in Section 7.1. The diagonal entries of $\Lambda$ range from 0.52 to 148.82 with a mean of 45.68. We set $T=10I_{d}$ and $\mu=0$ to represent an uninformative initialization. The stochastic gradients with respect to $\mu$ and $T$ are computed for each divergence by generating $z\sim{\text{N}}(0,I_{d})$ for 1000 iterations. The standard deviation of these gradient estimates are calculated for $\mu_{i}$ and $T_{ii}$ for $i=1,\dots,d$, and summarized using boxplots in Fig 4. The $y$-axis of the boxplots has a log scale. KLD has the smallest standard deviation, followed by SD, while the standard deviation of FD is much larger than SD and KLD for both $\mu$ and $T$. Although $\Lambda$ is not a diagonal matrix, these findings are consistent with our earlier analysis. This example highlights the difficulty in using SGD to minimize the FD and SD due to the much larger variance in gradient estimates relative to KLD, which motivates an alternative optimization procedure described next.
6 SGD based on batch approximation
FDb SDb Initialize $\mu$ and $T$. For $t=1,\dots N$, 1. Generate $z_{i}\sim{\text{N}}(0,I_{d})$ and compute $\theta_{i}=\mu+T^{-\top}z_{i}$ for $i=1,\dots B$. 2. Compute $g_{h}(\theta_{i})$ for $i=1,\dots B$ and the summary statistics in (8). 3. $U=C_{\theta}+(\mu-\overline{\theta})(\mu-\overline{\theta})^{\top}$. 4. $g_{\mu}=2TT^{\top}(\mu-\overline{\theta})-2\overline{g}_{h}$. 5. $\mu\leftarrow\mu-\rho_{t}TT^{\top}g_{\mu}$ 6. $W=C_{\theta g}-(\mu-\overline{\theta})\overline{g}_{h}^{\top}$. 7. $g_{T}=2(W+W^{\top}+TT^{\top}U+UTT^{\top})T$. 5. $\mu\leftarrow\mu-\rho_{t}g_{\mu}$. 6. $V=C_{g}+\overline{g}_{h}\overline{g}_{h}^{\top}$. 7. $g_{T}=2(UT-T^{-\top}T^{-1}VT^{-\top})$. 8. $T^{*}\leftarrow T^{*}-\rho_{t}Dg_{T}$. 9. Obtain $T$ from $T^{*}$.
Table 5: SGD algorithms based on batch approximated FD and SD.
Unbiased gradient estimates based on the reparametrization trick involve the Hessian $\nabla_{\theta}^{2}\log h(\theta)$, which is computationally expensive, requires high storage and increases the variability in gradient estimates. This can lead to instability in minimizing the FD and SD. To address these challenges, we propose an alternative approach that eliminates dependence on the Hessian, by minimizing biased estimates of the FD and SD computed using a batch of samples randomly selected from the current variational approximation at each iteration. First, the SD and FD can be written respectively as
$S_{q_{\lambda}}(\lambda)$ $={\text{E}}_{q_{\lambda}}\|g_{h}(\theta)+\Sigma^{-1}(\theta-\mu)\|^{2}_{\Sigma}$
$={\text{E}}_{q_{\lambda}}\left\{g_{h}(\theta)^{\top}\Sigma g_{h}(\theta)+2g_{h}(\theta)^{\top}(\theta-\mu)\right.$
$\left.\quad+(\theta-\mu)^{\top}\Sigma^{-1}(\theta-\mu)\right\},$
$F_{q_{\lambda}}(\lambda)$ $={\text{E}}_{q_{\lambda}}\|g_{h}(\theta)+\Sigma^{-1}(\theta-\mu)\|^{2}$
$={\text{E}}_{q_{\lambda}}\left\{g_{h}(\theta)^{\top}g_{h}(\theta)+2g_{h}(\theta)^{\top}\Sigma^{-1}(\theta-\mu)\right.$
$\left.\quad+(\theta-\mu)^{\top}\Sigma^{-2}(\theta-\mu)\right\},$
where $g_{h}(\theta)=\nabla_{\theta}\log h(\theta)$ and the subscript $q_{\lambda}$ emphasizes that expectation is with respect to $q_{\lambda}(\theta)$. To estimate SD and FD at the $t$-iteration, we generate $B$ samples $\{\theta_{1},\dots,\theta_{B}\}$ from the current estimate of the variational density $q_{t}(\theta)={\text{N}}(\theta|\mu^{(t)},\Sigma^{(t)})$. This can be done by generating $z_{i}\sim{\text{N}}(0,I_{d})$ and computing $\theta_{i}=\mu^{(t)}+T^{(t)-\top}z_{i}$ for $i=1,\dots,B$, where $\Sigma^{(t)}=T^{(t)-\top}T^{(t)-1}$. Let
$\begin{gathered}\overline{\theta}=\frac{1}{B}\sum_{i=1}^{B}\theta_{i},\quad\overline{g}_{h}=\frac{1}{B}\sum_{i=1}^{B}g_{h}(\theta_{i}),\quad\\ C_{\theta g}=\frac{1}{B}\sum_{i=1}^{B}(\theta_{i}-\overline{\theta})(g_{h}(\theta_{i})-\overline{g}_{h})^{\top},\\ C_{\theta}=\frac{1}{B}\sum_{i=1}^{B}(\theta_{i}-\overline{\theta})(\theta_{i}-\overline{\theta})^{\top},\quad\\ C_{g}=\frac{1}{B}\sum_{i=1}^{B}(g_{h}(\theta_{i})-\overline{g}_{h})(g_{h}(\theta_{i})-\overline{g}_{h})^{\top},\end{gathered}$ (8)
be summary statistics based on the batch of samples. Estimates of SD and FD are
$\hat{S}_{q_{t}}(\lambda)$ $=\frac{1}{B}\sum_{i=1}^{B}\left\{g_{h}(\theta_{i})^{\top}\Sigma g_{h}(\theta_{i})+2g_{h}(\theta_{i})^{\top}(\theta_{i}-\mu)\right.$
$\left.\quad+(\theta_{i}-\mu)^{\top}\Sigma^{-1}(\theta_{i}-\mu)\right\}$
$=\text{\rm tr}(V\Sigma)+\text{\rm tr}(U\Sigma^{-1})+2\text{\rm tr}(W),$
$\hat{F}_{q_{t}}(\lambda)$ $=\frac{1}{B}\sum_{i=1}^{B}\left\{g_{h}(\theta_{i})^{\top}g_{h}(\theta_{i})+2g_{h}(\theta_{i})^{\top}\Sigma^{-1}(\theta_{i}-\mu)\right.$
$\left.\quad+(\theta_{i}-\mu)^{\top}\Sigma^{-2}(\theta_{i}-\mu)\right\}$
$=\text{\rm tr}(V)+\text{\rm tr}(U\Sigma^{-2})+2\text{\rm tr}(W\Sigma^{-1}),$
where $U=C_{\theta}+(\mu-\overline{\theta})(\mu-\overline{\theta})^{\top}$, $V=C_{g}+\overline{g}_{h}\overline{g}_{h}^{\top}$ and $W=C_{\theta g}-(\mu-\overline{\theta})\overline{g}_{h}^{\top}$. Differentiating with respect to $\mu$ and $T$,
$\nabla_{\mu}\hat{S}_{q_{t}}(\lambda)$ $=2\Sigma^{-1}(\mu-\overline{\theta})-2\overline{g}_{h},$ (9)
$\nabla_{\mu}\hat{F}_{q_{t}}(\lambda)$ $=\Sigma^{-1}\nabla_{\mu}\hat{S}_{q_{t}}(\lambda),$
$\nabla_{\text{\rm vech}(T)}\hat{S}_{q_{t}}(\lambda)$ $=2\text{\rm vech}(UT-\Sigma VT^{-\top}),$
$\nabla_{\text{\rm vech}(T)}\hat{F}_{q_{t}}(\lambda)$ $=2\text{\rm vech}\{(W+W^{\top}+\Sigma^{-1}U$
$\quad+U\Sigma^{-1})T\}.$
These gradient estimates of SD and FD are biased because the $\theta$’s are replaced by samples $\{\theta_{1},\dots,\theta_{B}\}$ generated from $q_{t}(\theta)={\text{N}}(\theta|\mu^{(t)},\Sigma^{(t)})$, and are no longer functions of $(\mu,\Sigma)$ when we derive the gradients. The dependence on $\lambda$ through $\theta$ has been cut. On the other hand, the reparametrization trick in Section 5 produces unbiased estimates because the $\theta$’s are regarded as samples from $q(\theta)={\text{N}}(\theta|\mu,\Sigma)$, and remain functions of $(\mu,\Sigma)$ when the chain rule is applied to find the gradients.
With the batch approximation, all gradients are independent of the Hessian, which reduces computation costs significantly and enhances stability during optimization. As before, we only update elements of $\text{\rm vech}(T)$ not fixed by sparsity, and ensure positivity of diagonal entries in $T$ by applying a transformation. SGD algorithms for updating $(\mu,T)$ based on minimizing the batch approximated FD and SD are outlined in Table 5.
6.1 Interpretation and related methods
Previously, Elkhalil et al. (2021) have designed autoencoders based on minimizing a batch approximation of the Fisher divergence using SGD. Cai et al. (2024) also proposed a BaM algorithm that derived closed form updates of $(\mu,\Sigma)$ by minimizing the objective,
$\hat{S}_{q_{t}}(\lambda)+(2/\rho_{t})\text{\rm KL}(q_{t}\|q_{\lambda}),$
with respect to $\lambda$ at the $t$th iteration, where $\rho_{t}=Bd/t$ is the learning rate. BaM can be interpreted as a proximal point method that produces a sequence of variational densities $q_{0},q_{1},\dots$ such that $q_{t+1}$ matches the scores $g_{h}(\theta)$ at $\{\theta_{1},\dots,\theta_{B}\}$ on average better than $q_{t}$, while the KLD based penalty term ensures stability by preventing $q_{t+1}$ from deviating too much from $q_{t}$. Similarly, the SGD algorithm in Table 5 can be interpreted as minimizing
$\hat{S}_{q_{t}}(\lambda)+(1/2\rho_{t})\left\|\lambda-\lambda_{t}\right\|^{2}$
with respect to $\lambda$, where an $\ell_{2}$ penalty is used instead, and a linear approximation of $\hat{S}_{q_{t}}(\lambda)$ at $\lambda_{t}$ is considered. Then
$\lambda_{t+1}$ $=\arg\min_{\lambda}\left\{\hat{S}_{q_{t}}(\lambda_{t})+\nabla_{\lambda}\hat{S}_{q_{t}}(\lambda_{t})^{\top}(\lambda-\lambda_{t})\right.$
$\quad\left.+\ (1/2\rho_{t})\left\|\lambda-\lambda_{t}\right\|^{2}\right\}=\lambda_{t}-\rho_{t}\nabla_{\lambda}\hat{S}_{q_{t}}(\lambda_{t}),$
which corresponds to the SGD update with step size $\rho_{t}$ employed in Table 5. This discussion extends similarly to the Fisher divergence.
Figure 5: Progression of SDb for a Gaussian target where the red *’s mark the randomly chosen samples.
Instead of viewing (9) as biased estimates of the gradients of SD and FD, we can consider $\hat{S}_{q_{t}}(\lambda)$ and $\hat{F}_{q_{t}}(\lambda)$ as new objectives, which measure the divergence between $q_{\lambda}(\theta)$ and $p(\theta|y)$ based on their gradients evaluated at randomly selected samples at each iteration. Indeed, $\hat{S}_{q_{t}}(\lambda)$ and $\hat{F}_{q_{t}}(\lambda)$ reduce to zero when $q_{\lambda}(\theta)=p(\theta|y)$. At each iteration $t$, $q_{t+1}$ updates $q_{t}$ so as to reduce the difference in gradients between $q_{t}$ and the true posterior when evaluated on the randomly selected batch of samples. As $q_{t}$ converges to $p(\theta|y)$, the samples also shift towards the region where the true posterior has high probability mass. Figure 5 illustrates the progression of SDb with batch size $B=5$ using ${\text{N}}(2,1.5)$ as target. As optimization proceeds, SDb gradually refines the variational density by emulating the gradients of the target on the batch of samples and converges steadily toward the target.
The SGD algorithm in Table 5 differs from BaM in several key aspects. While BaM relies on KLD regularization to stabilize updates and has closed-form updates for $(\mu,\Sigma)$, we use an $\ell_{2}$ penalty and a linearization of the batch approximation leading to SGD. By avoiding linearization of $\hat{S}_{q_{t}}(\lambda)$ and SGD, the number of iterations required for convergence is reduced in BaM, but each BaM iteration is expensive as the closed form update of $\Sigma$ involves inverting a $d\times d$ matrix with cost of $\mathcal{O}(d^{3})$, although this can be reduced to $\mathcal{O}(d^{2}B+B^{3})$ for small batch size $B$ through low rank solvers. This high cost can result in long runtimes in high dimensions. Moreover, BaM is designed for full covariance matrices and it is not clear how sparsity can be enforced in the precision matrix to take advantage of the posterior conditional independence structure in hierarchical models. BaM can also run into instability and numerical issues with ill-conditioned matrices in practice, which may not be alleviated even with larger batch sizes. On the other hand, SGD allows updating of the Cholesky factor of the precision matrix, where sparse structures can be easily enforced. Smaller batches can also be used, which further reduces the computation and storage burden. While BaM is suited for full covariance Gaussian VI, our approach provides a more scalable and stable alternative for high-dimensional hierarchical models with conditional independence structure.
6.2 Batch approximated objective under mean-field
Next, we investigate behavior of the batch approximated FD and SD under the mean-field assumption considered in Section 2. Suppose the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$ with non-diagonal precision matrix $\Lambda$, and the variational approximation $q(\theta)$ is ${\text{N}}(\mu,\Sigma)$ where $\Sigma$ is a diagonal matrix. Using $B>1$ samples $\{\theta_{1},\dots,\theta_{B}\}$ from an estimate of $q$, $\hat{q}(\theta)={\text{N}}(\hat{\mu},\hat{\Sigma})$, where $\hat{\Sigma}$ is also a diagonal matrix, the batch approximated SD and FD are
$\hat{S}_{q}(\lambda)$ $=\sum_{i=1}^{d}(V_{ii}\Sigma_{ii}+U_{ii}\Sigma^{-1}_{ii})+2\text{\rm tr}(W),$
$\hat{F}_{q}(\lambda)$ $=\sum_{i=1}^{d}(U_{ii}\Sigma^{-2}_{ii}+2W_{ii}\Sigma^{-1}_{ii})+\text{\rm tr}(V).$
Lemma 2.
$\hat{S}_{q}(\lambda)$ is minimized at $\Sigma^{\hat{S}}_{ii}=\sqrt{C_{\theta,ii}/C_{g,ii}}$ and $\mu_{i}^{\hat{S}}=\overline{\theta}_{i}+\overline{g}_{h,i}\Sigma^{\hat{S}}_{ii}$ for $i=1,\dots,d$. If the diagonal entries of $C_{\theta g}$ are all negative, then $\hat{F}_{q}(\lambda)$ is minimized at $\Sigma^{\hat{F}}_{ii}=-C_{\theta,ii}/C_{\theta g,ii}$ and $\mu_{i}^{\hat{F}}=\overline{\theta}_{i}+\overline{g}_{h,i}\Sigma^{\hat{F}}_{ii}$ for $i=1,\dots,d$.
Proof.
Differentiating $\hat{S}_{q}(\lambda)$ and $\hat{F}_{q}(\lambda)$ with respect to $\mu$ and $\Sigma_{ii}$, we obtain
$\nabla_{\mu}\hat{S}_{q}(\lambda)$ $=2\Sigma^{-1}(\mu-\overline{\theta})-2\overline{g}_{h},$
$\nabla_{\mu}\hat{F}_{q}(\lambda)$ $=\Sigma^{-1}\nabla_{\mu}\hat{S}_{q}(\lambda),$
$\nabla_{\Sigma_{ii}}\hat{S}_{q}(\lambda)$ $=V_{ii}-U_{ii}\Sigma_{ii}^{-2},$
$\nabla_{\Sigma_{ii}}\hat{F}_{q}(\lambda)$ $=-2\Sigma_{ii}^{-2}(U_{ii}\Sigma_{ii}^{-1}+W_{ii}).$
Setting these derivatives to zero yields
[…]
Solving these equations simultaneously, we obtain
$V_{ii}(\Sigma^{\hat{S}}_{ii})^{2}$ $=C_{\theta,ii}+(\Sigma^{\hat{S}}_{ii})^{2}\overline{g}_{h,i}^{2}\implies\Sigma^{\hat{S}}_{ii}=\sqrt{C_{\theta,ii}/C_{g,ii}},$
$\Sigma^{\hat{F}}_{ii}$ =$=-\{C_{\theta,ii}+(\Sigma_{ii}^{\hat{F}})^{2}\overline{g}_{h,i}^{2}\}/\{C_{\theta g,ii}-\Sigma_{ii}^{\hat{F}}\overline{g}_{h,i}^{2}\}$
$\quad\implies\Sigma^{\hat{F}}_{ii}=-C_{\theta,ii}/C_{\theta g,ii}.$
Plugging these values into (10) yields corresponding values for $\mu_{i}^{\hat{S}}$ and $\mu_{i}^{\hat{F}}$. ∎
Next, we study limiting behavior of the batch approximated SD and FD as the batch size $B\to\infty$. Theorem 2 relies on the limits of summary statistics in (8) presented in Lemma 3.
Lemma 3.
Suppose $\{\theta_{1},\dots,\theta_{B}\}$ are samples from $\hat{q}(\theta)={\text{N}}(\hat{\mu},\hat{\Sigma})$ and the target is $p(\theta|y)={\text{N}}(\nu,\Lambda^{-1})$. As $B\rightarrow\infty$, the summary statistics in (8) satisfy
[…]
Proof.
The first two results follow directly from the law of large numbers. For the target, $g_{h}(\theta_{i})=-\Lambda(\theta_{i}-\nu)$. Thus $\overline{g}_{h}=-\Lambda(\overline{\theta}-\nu)$ and $g_{h}(\theta_{i})-\overline{g}_{h}=-\Lambda(\theta_{i}-\overline{\theta})$.
$\therefore\quad C_{g}$ $=\frac{1}{B}\sum_{i=1}^{B}\Lambda(\theta_{i}-\overline{\theta})(\theta_{i}-\overline{\theta})^{\top}\Lambda=\Lambda C_{\theta}\Lambda,$
$C_{\theta g}$ $=-\frac{1}{B}\sum_{i=1}^{B}(\theta_{i}-\overline{\theta})(\theta_{i}-\overline{\theta})^{\top}\Lambda=-C_{\theta}\Lambda.$
[…]
Theorem 2.
Suppose the target $p(\theta|y)$ is ${\text{N}}(\nu,\Lambda^{-1})$. Let the variational approximation $q(\theta)$ be ${\text{N}}(\mu,\Sigma)$, and $\hat{q}(\theta)={\text{N}}(\theta|\hat{\mu},\hat{\Sigma})$ be an estimate of $q(\theta)$, where $\Sigma$ and $\hat{\Sigma}$ are both diagonal matrices. As the batch size $B\rightarrow\infty$, the batch approximated SD and FD are minimized at $(\mu^{\hat{S}},\Sigma^{\hat{S}})$ and $(\mu^{\hat{F}},\Sigma^{\hat{F}})$ respectively, where
[…]
Proof.
Results can be obtained by applying the continuous mapping theorem on Lemma 2 and using the results in Lemma 3. Note that $(\Lambda\hat{\Sigma}\Lambda)_{ii}=\sum_{j=1}^{d}\hat{\Sigma}_{jj}\Lambda_{ij}^{2}$ and $(\hat{\Sigma}\Lambda)_{ii}=\hat{\Sigma}_{ii}\Lambda_{ii}$. ∎
[…]
7 Applications
We evaluate the performance of algorithms in Tables 4 and 5 by applying them to logistic regression, GLMMs and stochastic volatility models, and compare their results with BaM and MCMC. MCMC sampling is performed using RStan by running two chains in parallel, each with 50,000 iterations. The first half is discarded as burn-in, while the remaining 50,000 draws are used to compute kernel density estimates, regarded as the gold standard.
To evaluate the multivariate accuracy of variational approximation relative to MCMC, we use maximum mean discrepancy (MMD, Zhou et al., 2023). We calculate $M^{*}=-\log(\text{MMD}^{2}_{u}+10^{-5})$, where
$\text{MMD}^{2}_{u}$ $=\frac{1}{m(m-1)}\sum_{i\neq j}^{m}[k(\textbf{x}_{v}^{(i)},\textbf{x}_{v}^{(j)})+k(\textbf{x}_{g}^{(i)},\textbf{x}_{g}^{(j)})$
$\quad-k(\textbf{x}_{v}^{(i)},\textbf{x}_{g}^{(j)})-k(\textbf{x}_{v}^{(j)},\textbf{x}_{g}^{(i)})],$
$\textbf{x}_{v}^{(1)},\dots,\textbf{x}_{v}^{(m)}$ and $\textbf{x}_{g}^{(1)},\dots,\textbf{x}_{g}^{(m)}$ represent independent samples drawn from the variational approximation and MCMC respectively, $k$ is the radial basis kernel function and $m=1000$. $M^{*}$ is computed 50 times for each variational approximation and a higher value indicates better multivariate accuracy. In addition, we assess the ability to capture the marginal mode of each variable accurately using the normalized absolute difference in mode ($|\mu-\mu^{*}|/\sigma^{*}$), where $\mu^{*}$ and $\sigma^{*}$ denote the mode and standard deviation of each variable based on MCMC samples.
To assess convergence, we track unbiased estimates of the lower bound, $\hat{{\mathcal{L}}}$, averaged over every 1000 iterations for SGD methods and 50 iterations for BaM, to reduce noise. Less iterations are used for averaging in BaM, as it uses closed form updates, which leads to more stable trajectories with reduced oscillations. Moreover, BaM usually requires a larger batch size and converges faster than SGD methods. Each algorithm is terminated when the gradient of a linear regression line fitted to the past five lower bound averages is less than 0.01, or when the maximum number of iterations is reached. The batch size of FDb and SDb is adjusted according to model complexity, and larger batch sizes are used for more complex models. All experiments are performed on a 16GB Apple M1 computer, using R and Julia 1.11.2.
7.1 Logistic regression
KLD FDr SDr FDb SDb BaM MCMC
MMD German 8.82$\pm$1.72 1.11$\pm$0.02 0.75$\pm$0.01 2.86$\pm$0.07 8.37$\pm$1.28 8.29$\pm$1.35 -
a4a 2.49$\pm$0.03 0.71$\pm$0.02 0.71$\pm$0.01 1.61$\pm$0.02 3.04$\pm$0.04 5.22$\pm$0.08 -
$\frac{|\mu-\mu^{*}|}{\sigma^{*}}$ German 0.08$\pm$0.05 0.55$\pm$0.68 0.74$\pm$0.62 0.25$\pm$0.27 0.08$\pm$0.06 0.08$\pm$0.05 -
a4a 0.18$\pm$0.15 0.45$\pm$0.52 0.38$\pm$0.43 0.40$\pm$0.43 0.15$\pm$0.15 0.14$\pm$0.16 -
time German 3.2 (35) 15.0 (60) 9.8 (39) 6.2 (26) 3.8 (15) 1.6 (0.35) 354.4
a4a 19.2 (44) 108.5 (31) 55.2 (16) 14.0 (10) 70.4 (48) 13.6 (1.05) 28556.1
Table 6: Logistic regression. Estimates of mean and standard deviation for MMD and normalized absolute difference in mode.
Runtime is in seconds and number of iterations (in thousands) is given in brackets.
Consider a logistic regression model where $y=(y_{1},\dots,y_{n})^{\top}$ represents $n$ independent binary responses. Each $y_{i}$ follows a Bernoulli distribution with success probability $p_{i}$, modeled as
$\text{\rm logit}(p_{i})=X_{i}^{\top}\theta\quad\text{for}\quad i=1,\dots,n,$
where $X_{i}\in\mathbb{R}^{d}$ denotes the covariates of the $i$th observation and $\theta\in\mathbb{R}^{d}$ denotes the unknown coefficients, which is assigned the prior ${\text{N}}(0,\sigma_{0}^{2}I_{d})$ with $\sigma_{0}^{2}=100$. Here, the precision matrix of the Gaussian variational approximation is not sparse, but a full matrix. The log joint density of the model, gradient and Hessian are given by
$\log h(\theta)$ $=y^{\top}X\theta-\sum_{i=1}^{n}\log\{1+\exp(X_{i}^{\top}\theta)\}$
$\quad-\tfrac{d}{2}\log(2\pi\sigma_{0}^{2})-\theta^{\top}\theta/(2\sigma_{0}^{2}),$
$\nabla_{\theta}\log h(\theta)$ $=X^{\top}(y-w)-\theta/\sigma_{0}^{2},$
$\nabla_{\theta}^{2}\log h(\theta)$ $=-X^{\top}WX-I_{d}/\sigma_{0}^{2},$
where $w=(w_{1},\dots,w_{n})^{\top}$, $w_{i}=\{1+\exp(-X_{i}^{\top}\theta)\}^{-1}$ for $i=1,\dots,n$, $W$ is an $n\times n$ diagonal matrix with diagonal entries $w_{i}(1-w_{i})$ and $X=(X_{1},\dots,X_{n})^{\top}$.
We fit the logistic regression model to two real datasets from the UCI machine learning repository. The first is the German credit data, which consists of $1000$ individuals classified as having a “good” or “bad” credit risk, and $20$ attributes. All quantitative predictors are standardized to have mean zero and standard deviation one, while qualitative predictors are encoded using dummy variables. The second is the Adult data with $48,842$ observations, which is used to predict whether an individual’s annual income exceeds \$$50,000$ based on $14$ attributes. For MCMC to be feasible, we use the preprocessed a4a data at www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/binary.html, which has $4781$ training samples derived from the Adult data. After preprocessing, $d=49$ for German credit data and $d=124$ for a4a data. As the a4a data is high dimensional and has a large number of observations, we only generate $10,000$ MCMC samples from two parallel chains, each consisting of $10,000$ iterations. We use a batch size of $B=3$ for FDb and SDb, and $B=50$ for BaM. The maximum number of iterations is $60,000$.
Figure 6: German credit and a4a data. Average lower bound over every 1000 iterations.
Fig 6 shows progression of the lower bound for SGD methods. FDr and SDr converge very slowly, and attain much poorer lower bounds than other methods at termination. This is likely due to the high variance in their gradient estimates, as discussed in Section 5.1. In contrast, SDb converges rapidly and achieves the highest lower bound within the first one thousand iterations, surpassing even KLD. The lower bound achieved by FDb, on the other hand, is lower than KLD and SDb although it performs better than FDr and SDr.
From the MMD results in Table 6, FDr and SDr produce much poorer variational approximations than KLD, while FDb and SDb provide significant improvements over FDr and SDr. In particular, SDb produced results comparable to KLD for the German credit data, and better than KLD for the a4a data. In terms of runtime, KLD is the fastest among SGD methods. For the German credit data, FDr, SDr, FDb and SDb each takes $\sim$0.25s per 1000 iterations, but FDb and SDb require less iterations to converge. For the a4a data, FDr and SDr require $\sim$3.5s per 1000 iterations compared to $\sim$1.5s for FDb and SDb, due to the larger cost of computing the Hessian in high dimensions. BaM converges most rapidly, outperforming all SGD methods in runtime. Its MMD values are comparable to KLD for the German credit data, and much higher than SGD methods for the a4a data.
Overall, MMD values for optimizing FD and SD based on batch approximation are consistently higher than those based on the reparameterization trick. In terms of capturing the marginal mode accurately, SDb and BaM are comparable to KLD for the German credit data, and better than KLD for the more challenging a4a data. All variational approximations are at least an order of magnitude faster than MCMC, and the speedup is more apparent as the dimension and number of observations increase.
7.2 Generalized linear mixed models
Let $y_{i}=(y_{i1},\dots,y_{in_{i}})^{\top}$ denote the $n_{i}$ observations for the $i$th subject and $y=(y_{1}^{\top},\dots,y_{n}^{\top})^{\top}$. Each $y_{ij}$ is distributed according to a density in the exponential family, and a smooth invertible link function $g(\cdot)$ relates its mean $\mu_{ij}$ to a linear predictor $\eta_{ij}$ such that
$g(\mu_{ij})=\eta_{ij}=X_{ij}^{\top}\beta+Z_{ij}^{\top}b_{i}$
for $i=1,\dots,n$, $j=1,\dots,n_{i}$. Here, $\beta\in\mathbb{R}^{p}$ is the fixed effect, $b_{i}\in\mathbb{R}^{r}$ is the random effect characterizing the $i$th subject, and $X_{ij}\in\mathbb{R}^{p}$ and $Z_{ij}\in\mathbb{R}^{r}$ are the covariates. We assume $b_{i}\sim{\text{N}}(0,G^{-1})$ and let $G=WW^{\top}$ be the Cholesky decomposition of precision matrix $G$, where $W$ is a lower triangular matrix with positive diagonal entries. For unconstrained optimization of $W$, we introduce $W^{*}$ such that $W_{ii}^{*}=\log(W_{ii})$ and $W_{ij}^{*}=W_{ij}$ if $i\neq j$, and let $\zeta=\text{vech}(W^{*})$. Normal priors, $\beta\sim{\text{N}}(0,\sigma_{0}^{2}I_{p})$ and $\zeta\sim N(0,\sigma_{0}^{2}I_{r(r+1)/2})$, where $\sigma_{0}^{2}=100$ are assigned. The global variables are $\theta_{G}=(\beta^{\top},\zeta^{\top})^{\top}$ and the local variables are $\theta_{L}=(b_{1}^{\top},\dots,b_{n}^{\top})^{\top}$. Focusing on GLMMs with canonical links and the one-parameter exponential family, the log joint density of the model can be written as
$\log h(\theta)$ $=\sum_{i=1}^{n}\sum_{j=1}^{n_{i}}\log p(y_{ij}|\beta,b_{i})+\sum_{i=1}^{n}\log p(b_{i}|\zeta)$
$\quad+\log p(\beta)+\log p(\zeta)$
$=\sum_{i,j}\{y_{ij}\eta_{ij}-A(\eta_{ij})\}+n\log|W|$
$\quad-\frac{1}{2}\sum_{i=1}^{n}b_{i}^{\top}WW^{\top}b_{i}-\frac{\beta^{\top}\beta}{2\sigma_{\beta}^{2}}-\frac{\zeta^{\top}\zeta}{2\sigma_{\zeta}^{2}}+C,$
where $A(\cdot)$ is the log-partition function and $C$ is a constant independent of $\theta$. For instance, $A(x)=\log(1+e^{x})$ for Bernoulli-distributed binary responses and $A(x)=\exp(x)$ for Poisson-distributed count responses. Let $H=\nabla_{\theta}^{2}\log h(\theta)$ and $H_{\theta_{i},\theta_{j}}=\nabla_{\theta_{i},\theta_{j}}^{2}\log h(\theta)$. The Hessian,
$H=\begin{bmatrix}H_{b_{1},b_{1}}&\ldots&0&H_{b_{1},\theta_{G}}\\ \vdots&\ddots&\vdots&\vdots\\ 0&\dots&H_{b_{n},b_{n}}&H_{b_{n},\theta_{G}}\\ H_{\theta_{G},b_{1}}&\ldots&H_{\theta_{G},b_{n}}&H_{\theta_{G}}\\ \end{bmatrix},$
has a sparse structure similar to $\Omega$. Both $\nabla_{\theta}h(\theta)$ and $H$ are derived in the supplement.
First, consider the epilepsy data (Thall and Vail, 1990) from a clinical trial with $n=59$ patients, who were randomly assigned to a drug, progabide ($\text{Trt}=1$), or a placebo ($\text{Trt}=0$). The response is the number of seizures experienced by each patient during four follow-up periods. Covariates include logarithm of the patient’s age at baseline, which is centered by subtracting the mean (Age), logarithm of 1/4 the number of seizures prior to the trial (Base), visit number coded as --0.3, --0.1, 0.1, 0.3 (Visit), and an indicator of the 4th visit (V4). We consider Poisson mixed models with random intercepts and slopes (Breslow and Clayton, 1993),
$\text{Epi I:}\log\mu_{ij}=\beta_{0}+\beta_{\text{Base}}\text{Base}_{i}+\beta_{\text{Trt}}\text{Trt}_{i}+\beta_{\text{Age}}\text{Age}_{i}\\ +\beta_{\text{Base}\text{Trt}}\text{Base}_{i}\text{Trt}_{i}+\beta_{\text{V4}}\text{V4}_{ij}+b_{i},\\ \text{Epi II:}\log\mu_{ij}=\beta_{0}+\beta_{\text{Base}}\text{Base}_{i}+\beta_{\text{Trt}}\text{Trt}_{i}+\beta_{\text{Age}}\text{Age}_{i}\\ +\beta_{\text{Base}\text{Trt}}\text{Base}_{i}\text{Trt}_{i}+\beta_{\text{Visit}}\text{Visit}_{ij}+b_{i1}+b_{i2}\text{Visit}_{ij},$
for $i=1,...,n$, $j=1,...,4$.
KLD FDr SDr FDb SDb BaM MCMC
MMD Epi I 6.37$\pm$0.19 0.04$\pm$0.02 -0.06$\pm$0.02 2.89$\pm$0.05 6.48$\pm$0.22 9.51$\pm$1.52 -
Epi II 6.34$\pm$0.13 0.04$\pm$0.03 0.06$\pm$0.03 4.65$\pm$0.05 5.97$\pm$0.17 6.41$\pm$0.22 -
Toenail 4.17$\pm$0.05 0.35$\pm$0.01 0.11$\pm$0.01 1.42$\pm$0.01 2.53$\pm$0.02 2.46$\pm$0.02 -
Polypharmacy 5.66$\pm$0.05 0.49$\pm$0.00 0.19$\pm$0.01 2.40$\pm$0.01 3.79$\pm$0.03 3.80$\pm$0.03 -
$\frac{|\mu-\mu^{*}|}{\sigma^{*}}$ Epi I 0.07$\pm$0.05 2.04$\pm$1.62 2.30$\pm$2.09 0.28$\pm$0.19 0.07$\pm$0.05 0.07$\pm$0.05 -
Epi II 0.09$\pm$0.07 2.14$\pm$1.79 2.55$\pm$2.40 0.19$\pm$0.16 0.09$\pm$0.08 0.09$\pm$0.07 -
Toenail 0.21$\pm$0.14 1.20$\pm$1.30 1.57$\pm$2.17 0.68$\pm$0.65 0.36$\pm$0.20 0.38$\pm$0.22 -
Polypharmacy 0.18$\pm$0.10 1.01$\pm$1.19 1.34$\pm$2.12 0.44$\pm$0.38 0.22$\pm$0.13 0.23$\pm$0.13 -
$\frac{\sigma}{\sigma^{*}}$ Epi I 0.95$\pm$0.06 0.76$\pm$0.26 0.72$\pm$0.34 0.81$\pm$0.21 0.94$\pm$0.05 0.99$\pm$0.02 -
Epi II 0.94$\pm$0.08 0.94$\pm$0.29 0.69$\pm$0.28 0.88$\pm$0.18 0.95$\pm$0.08 0.97$\pm$0.07 -
Toenail 0.87$\pm$0.05 0.35$\pm$0.14 0.10$\pm$0.05 0.66$\pm$0.13 0.76$\pm$0.11 0.75$\pm$0.12 -
Polypharmacy 0.94$\pm$0.03 0.39$\pm$0.12 0.13$\pm$0.05 0.80$\pm$0.09 0.86$\pm$0.07 0.86$\pm$0.07 -
time Epi I 2.3 (40) 16.7 (47) 4.3 (12) 6.6 (24) 8.9 (26) 0.8 (0.3) 95.9
Epi II 3.9 (37) 73.3 (60) 26.0 (21) 16.7 (25) 28.5 (31) 13.1 (1.55) 380.8
Toenail 2.8 (11) 117.4 (30) 27.7 (7) 93.7 (30) 35.0 (10) 32.5 (0.3) 171.1
Polypharmacy 5.1 (10) 346.1 (30) 94.7 (7) 285.8 (30) 131.8 (11) 112.5 (0.4) 506.2
Table 7: GLMMs: Estimates of mean and standard deviation for MMD, normalized absolute difference in mode and standard deviation ratio.
Runtime is in seconds and number of iterations (in thousands) is given in brackets.
Next, consider the toenail data (De Backer et al., 1998) from a clinical trial comparing two oral antifungal treatments for toenail infections. Each of 294 patients was evaluated for up to seven visits, resulting in a total of 1908 observations. Patients were randomized to receive 250 mg of terbinafine ($\text{Trt}=1$) or 200 mg of itraconazole ($\text{Trt}=0$) per day. The response variable is binary, with 0 indicating no or mild nail separation and 1 for moderate or severe separation. Visit times in months ($t$) are standardized to have mean 0 and variance 1. A logistic random intercept model is fitted to this data,
$\text{logit}(\mu_{ij})=\beta_{0}+\beta_{\text{Trt}}\text{Trt}_{i}+\beta_{t}t_{ij}+\beta_{\text{Trt}\times t}\text{Trt}_{i}\times t_{ij}+b_{i},$
for $i=1,...,294,1\leq j\leq 7$.
[…]
logit $(\mu_{ij})=\beta_{0}+\beta_{\text{Gender}}\text{Gender}_{i}+\beta_{\text{Race}}\text{Race}_{i}$
$+\beta_{\text{Age}}\text{Age}_{ij}+\beta_{\text{MHV1}}\text{MHV1}_{ij}+\beta_{\text{MHV2}}\text{MHV2}_{ij}$
$+\beta_{\text{MHV3}}\text{MHV3}_{ij}+\beta_{\text{INPT}}\text{INPTMHV}_{ij}+b_{i},$
for $i=1,\dots,500$, $j=1,\dots,7$.
In this section, we use batch size $B=5$ for FDb and SDb. For BAM, $B=100$ for the epilepsy data and $B=1000$ for the toenail and polypharmacy data. A larger batch size is used in the latter cases as BAM is very prone to ill-conditioned updates and the convergence is very slow for smaller batch sizes. The maximum number of iterations for the epilepsy data is $60,000$, and reduced to $30,000$ for the toenail and polypharmacy data as they contain much more observations.
Figure 7: GLMMs: Average lower bound every 1000 iterations.
Fig 7 shows that SDb is among the fastest to converge among SGD methods, achieving a lower bound higher than KLD for Epi I, Epi II and polypharmacy, and comparable to KLD for toenail. While FDb converges rapidly for Epi I and Epi II, it fails to converge by the maximum iteration of 30,000 for toenail and polypharmacy. Overall, methods based on FD (FDr, FDb) seem to converge much more slowly than methods based on SD (SDr, SDb).
From the MMD results in Table 7, FDr and SDr have the lowest MMD. In contrast, FDb and SDb yield substantial improvements over their reparameterization trick-based counterparts, achieving much higher MMDs. Among SGD methods based on the weighted Fisher divergence, SDb has the highest MMD, even surpassing KLD for Epi I. BaM outperforms KLD for Epi I and Epi II, but its performance falters for toenail and polypharmacy, where the dimension of $\theta$ is higher, reflecting the challenges that BaM faces in high-dimensional cases.
KLD is able to capture the marginal posterior mode of each variable most accurately, with comparable performance from SDb and BaM. While BaM captured the marginal posterior variance most accurately for Epi I and Epi II, its performance falls behind KLD for the higher-dimensional toenail and polypharmacy data. SDr underestimates the marginal posterior variance most severely, which is reminiscent of the “variational collapse” problem it faces in the mean-field setting.
Figure 8: Marginal density estimates for some local and global variables in Epi I (first row) and polypharmacy (second row).
BaM is able to converge using the least number of iterations, by leveraging on closed-form updates and larger batch sizes. However, the computation cost per iteration of BaM is much higher than SGD methods, which can exploit the sparse structure of the precision matrix. This issue becomes more apparent as the dimension of $\theta$ increases. Among SGD methods, the runtime of KLD is the shortest. Methods based on FD tend to require more iterations to converge than those based on SD, resulting in longer runtime. SDb is able to converge in approximately the same number of iterations as KLD, although each iteration takes longer.
Fig 8 compares the marginal densities estimated using MCMC with variational approximations from KLD, SDb, and BaM for some variables in Epi I and polypharmacy. For Epi I in the first row, all variational approaches match the MCMC results very closely except for $\zeta$, where SDb underestimated the marginal posterior variance more severely than BaM and KLD. For the higher dimensional polypharmacy data in the second row, KLD matches the MCMC results most closely, while SDb and BaM display a higher tendency to underestimate the marginal posterior variance although the mode was captured more accurately in some cases.
7.3 Stochastic volatility models
The stochastic volatility model is widely used to capture the dynamic nature of financial time series. It provides an attractive alternative to constant volatility models like the Black-Scholes model (Black and Scholes, 1973), as the volatility of asset returns evolves over time according to a stochastic process. The response at time $t$ is
$y_{t}\sim{\text{N}}(0,\exp(\lambda+\sigma b_{t}))\quad\text{for}\quad t=1,\dots,n,$
where $\lambda\in\mathbb{R}$, $\sigma>0$, and the latent volatility process $b_{t}$ follows an autoregressive model of order one such that
$b_{t}$ $\sim{\text{N}}(\phi b_{t-1},1)\;\text{for}\;t=2,\dots,n,$
$b_{1}$ $\sim{\text{N}}(0,1/(1-\phi^{2})),$
where $\phi\in(0,1)$. To allow unconstrained updates, we apply the transformations, $\sigma=\exp(\alpha)$ and $\phi=\{\exp(-\psi)+1\}^{-1}$. The set of local variables are $\theta_{L}=(b_{1},\dots,b_{n})^{\top}$ and the set of global variables are $\theta_{G}=(\alpha,\lambda,\psi)^{\top}$. We consider the prior $\theta_{G}\sim{\text{N}}(0,\sigma_{0}^{2}I)$, where $\sigma_{0}^{2}=10$. For this model, $b_{i}$ is independent of $b_{j}$ given $\theta_{G}$ a posteriori if $|i-j|>1$ The log joint density is
$\log h(\theta)$ $=-\frac{n\lambda}{2}-\frac{\sigma}{2}\sum_{t=1}^{n}b_{t}-\frac{1}{2}\sum_{t=1}^{n}y_{t}^{2}\exp\{-\lambda-\sigma b_{t}\}$
$\quad-\frac{1}{2}\sum_{t=2}^{n}(b_{t}-\phi b_{t-1})^{2}+\frac{1}{2}\log(1-\phi^{2})$
$\quad-\frac{1}{2}b_{1}^{2}(1-\phi^{2})-\frac{\alpha^{2}}{2\sigma_{0}^{2}}-\frac{\lambda^{2}}{2\sigma_{0}^{2}}-\frac{\psi^{2}}{2\sigma_{0}^{2}}+C,$
where $C$ is a constant independent of $\theta$. The gradient $\nabla_{\theta}h(\theta)$ and the Hessian,
$\nabla_{\theta}^{2}\log h(\theta)=\begin{bmatrix}H_{b_{1},b_{1}}&H_{b_{1},b_{2}}&\ldots&0&H_{b_{1},\theta_{G}}\\ H_{b_{2},b_{1}}&H_{b_{2},b_{2}}&\ldots&0&H_{b_{2},\theta_{G}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dots&H_{b_{n},b_{n}}&H_{b_{n},\theta_{G}}\\ H_{\theta_{G},b_{1}}&H_{\theta_{G},b_{2}}&\ldots&H_{\theta_{G},b_{n}}&H_{\theta_{G},\theta_{G}}\\ \end{bmatrix},$
which has the same sparsity structure as $\Omega$ in the variational approximation, are derived in the supplement.
KLD FDr SDr FDb SDb BaM MCMC
MMD GBP 3.83$\pm$0.09 0.45$\pm$0.01 1.03$\pm$0.01 0.49$\pm$0.01 3.26$\pm$0.08 3.59$\pm$0.09 -
DEM 5.34 $\pm$ 0.27 0.36 $\pm$ 0.01 0.78 $\pm$ 0.01 0.40 $\pm$ 0.01 6.02 $\pm$ 0.15 - -
$\frac{|\mu-\mu^{*}|}{\sigma^{*}}$ GBP 0.07$\pm$0.05 0.43$\pm$0.43 0.36$\pm$0.32 0.38$\pm$0.37 0.04$\pm$0.03 0.05$\pm$0.04 -
DEM 0.09$\pm$0.07 1.15$\pm$0.91 1.19$\pm$0.74 0.99$\pm$0.77 0.06$\pm$0.05 - -
time GBP 7.1 (27) 12.7 (30) 6.6 (15) 621.8 (30) 419.2 (16) 1521.5 (3.25) 3573.2
DEM 9.7 (19) 27.3 (30) 23.7 (25) 2748.0 (30) 1281.3 (11) - 5764.9
Table 8: SSMs: Estimates of mean and standard deviation for MMD and normalized absolute difference in mode.
Runtime is in seconds and number of iterations (in thousands) is given in brackets.
We analyze two datasets from Garch in the R package Ecdat. The first contains $n=945$ observations of the weekday exchange rates of the U.S. Dollar against the British Pound (GBP) from 1 Oct 1981 to 28 June 1985. The second contains $n=1866$ observations of the weekday exchange rates for the U.S. Dollar against the German Deutschemark (DEM) from 2 Jan 1980 to 21 June 1987. For both datasets, the mean-corrected log-return series $\{y_{t}\}$ is derived from exchange rates $\{r_{t}\}$ using
$y_{t}=100\times\left\{\log\left(\frac{r_{t}}{r_{t-1}}\right)-\frac{1}{n}\sum_{i=1}^{n}\log\left(\frac{r_{i}}{r_{i-1}}\right)\right\}.$
We use $B=10$ for FDb and SDb, while BaM is executed with $B=50$ for GBP. We have tried larger batch sizes for BaM, but these resulted in severely ill-conditioned updates. For the higher-dimensional DEM, we cannot get BaM to converge despite trying various batch sizes. The challenge of inferring a full covariance matrix of order $d=1869$ for BaM here is immense, not to mention the computational cost of inverting matrices of this order. The maximum number of iterations is set as $30,000$.
From the MMD values in Table 8, FDr, SDr and FDb perform very poorly, compared to the rest. SDb yields the best performance among SGD methods optimizing the weighted Fisher divergence. Its MMD values are close to those of KLD for GBP, and even surpass KLD for the higher-dimensional DEM. SDb and BaM are also able to capture the mode more accurately than KLD for both datasets. In terms of runtime, KLD is the most efficient approach, attaining very high MMD values with the shortest runtime. While BaM requires fewer iterations to converge, the computational cost per iteration is much higher, leading to long runtimes. For instance, BaM requires 1521 seconds to converge for GBP, which is nearly triple the runtime of SDb. As the dimensionality increases, BaM also becomes more unstable and prone to ill-conditioned updates making convergence very difficult, while SDb provides a more reliable and computationally feasible approach for optimizing the batch approximated SD.
Figure 9: SSMs: Average lower bound over every 1000 iterations and boxplot of MMD values for different batch sizes for SDb.
Fig 9 presents the impact of varying the batch size for SDb in terms of the rate of convergence and quality of approximation measured by MMD. Increasing the batch size clearly leads to faster convergence and improved approximation quality. The total runtime as shown in the legends of the first row of Fig 9 is also reduced as lesser number of iterations is required for convergence. This suggests that larger batch sizes can enhance both the stability and accuracy of SDb.
Fig 10 shows the marginal posterior density estimates from MCMC, KLD and SDb ($B=10,100$) for some local variables and all the global variables from DEM. SDb can capture the marginal posterior mode more accurately than KLD, especially for each of the global variables, but has a higher tendency to underestimate the posterior variance. Increasing the batch size from 10 to 100 helps in reducing underestimation of the posterior variance.
Figure 10: Marginal density estimates for some local and global variables in DEM.
8 Conclusion
In this article, we evaluate the performance of Gaussian variational inference based on the weighted Fisher divergence by focusing on the FD and SD. First, we consider the mean-field assumption for both Gaussian and non-Gaussian targets. We demonstrate that FD and SD tend to underestimate the posterior variance more severely than KLD for Gaussian targets in this context. SD has the potential to capture the posterior mode more accurately than FD and KLD for skewed targets, but underestimates the variance more severely.
Next, we consider high-dimensional hierarchical models for which posterior conditional independence structure can be captured using a sparse precision matrix in the Gaussian variational approximation. To impose sparsity on the Cholesky factor of the precision matrix, we consider optimization based on SGD and propose two approaches based on the reparametrization trick and a batch approximation of the objective.
The reparametrization trick yields unbiased gradient estimates but involves a Hessian matrix, which can be computationally expensive and increases variability in the gradients, leading to reduced stability and slow convergence. To address these issues, we introduce an alternative that minimizes a biased estimate of the FD and SD using a random batch of samples at each iteration. This eliminates reliance on the Hessian matrix, improving stability. This approach can also be interpreted as optimizing a new objective that iteratively improves the match between gradients of the posterior and variational density, at sample points that shift gradually towards high probability regions of the posterior. We evaluate the behavior of this new objective under the mean-field assumption for Gaussian targets and show that it alleviates the variational collapse issue faced previously by SD.
The proposed methods are compared to KLD and BaM in applications involving logistic regression, GLMMs and stochastic volatility models. Extensive experiments reveal that FDr and SDr converge very slowly, often to suboptimal variational approximations. FDb and SDb provide substantial improvements over FDr and SDr, with SDb having superior performance in terms of convergence rate and quality of variational approximation attained. BaM, which relies on closed-form updates and hence requires fewer iterations to converge, is very effective for logistic regression. However, it is much less efficient than KLD for GLMMs and SSMs, and its performance gradually worsens as the dimension increases, eventually failing to converge. SDb has an advantage over BaM in high dimensions as it is able to impose sparsity on the precision matrix, remains feasible computationally and is more stable and less sensitive to poor initialization. SDb can also capture posterior modes more accurately than KLD although it is more prone to variance underestimation.
Several avenues for future research remain open. Our analysis focused on Gaussian variational approximations for FD and SD, and it will be valuable to investigate the performance of FD and SD under more flexible variational families. While we have used SGD for optimization, the choice of optimizer and associated hyperparameters significantly influences convergence behavior, and it will be useful to explore alternative optimization techniques that do not rely on SGD. Our findings also highlight the potential of the batch approximated SD, and it will be interesting to further investigate its properties in other contexts.
{funding}
Linda Tan’s research is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award MOE-T2EP20222-0002). David Nott’s research is supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 2 (MOE-T2EP20123-0009).
Supplementary Material
S1 Univariate non-Gaussian target
We begin by deriving some key expressions that will be used throughout our analysis of non-Gaussian target distributions. We assume the variational approximation $q(\theta)=\phi(\theta|\mu,\sigma^{2})$. Its log-density and gradient are
$\log q(\theta)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{2}\log(\sigma^{2})-\frac{(\theta-\mu)^{2}}{2\sigma^{2}},$
$\nabla_{\theta}\log q(\theta)$ $=-\frac{\theta-\mu}{\sigma^{2}}.$
The expectation of $\log q(\theta)$ with respect to $q(\theta)$ is
${\text{E}}_{q}\left\{\log q(\theta)\right\}=-\frac{1}{2}\log(2\pi\sigma^{2})-\frac{1}{2}.$
S1.1 Student’s $t$
The log-density and its gradient for $\theta\sim t(\nu)$ are
$\log p(\theta,y)$ $=\log\left\{\Gamma\left(\frac{\nu+1}{2}\right)\right\}-\frac{1}{2}\log(\pi\nu)$
$\quad-\log\left\{\Gamma\left(\frac{\nu}{2}\right)\right\}-\frac{\nu+1}{2}\log\left(1+\frac{\theta^{2}}{\nu}\right),$
$\nabla\log p(\theta,y)$ $=-(\nu+1)\frac{\theta}{\nu+\theta^{2}},$
where $\nu$ is the degree of freedom. The evidence lower bound for KLD is
${\mathcal{L}}$ $={\text{E}}_{q}\left\{\log p(\theta,y)-\log q(\theta)\right\}$
$=\log\left\{\Gamma\left(\frac{\nu+1}{2}\right)\right\}-\frac{1}{2}\log(\pi\nu)-\log\left\{\Gamma\left(\frac{\nu}{2}\right)\right\}$
$\quad-\frac{\nu+1}{2}{\text{E}}_{q}\log\left(1+\frac{\theta^{2}}{\nu}\right)+\frac{1}{2}\log(2\pi\sigma^{2})+\frac{1}{2}.$
The Fisher divergence is
$F(q\|p)$ $={\text{E}}_{q}\{\|\nabla_{\theta}\log p(\theta,y)-\nabla_{\theta}\log q(\theta)\|^{2}\}$
$={\text{E}}_{q}\left\{\left\|-(\nu+1)\frac{\theta}{\nu+\theta^{2}}+\frac{\theta-\mu}{\sigma^{2}}\right\|^{2}\right\}$
$=(\nu+1)^{2}{\text{E}}_{q}\left\{\frac{\theta^{2}}{(\nu+\theta^{2})^{2}}\right\}+{\text{E}}_{q}\left\{\frac{(\theta-\mu)^{2}}{\sigma^{4}}\right\}$
$\quad-2\frac{\nu+1}{\sigma^{2}}{\text{E}}_{q}\left\{\frac{\theta(\theta-\mu)}{\nu+\theta^{2}}\right\}$
$=(\nu+1)^{2}{\text{E}}_{q}\left\{\frac{\theta^{2}}{(\nu+\theta^{2})^{2}}\right\}+\frac{1}{\sigma^{2}}$
$\quad-\frac{2(\nu+1)}{\sigma^{2}}{\text{E}}_{q}\left\{\frac{\theta(\theta-\mu)}{\nu+\theta^{2}}\right\}.$
S1.2 Log transformed inverse Gamma
For the log transformed inverse gamma, let $a_{1}=a_{0}+n/2$ and $b_{1}=b_{0}+\sum_{i=1}^{n}y_{i}^{2}/2$. Then its log-density and gradient are given by
$\log p(y,\theta)$ $=-\frac{n}{2}\log(2\pi)+a_{0}\log b_{0}-\log\Gamma(a_{0})$
$\quad-a_{1}\theta-b_{1}\exp(-\theta),$
$\nabla_{\theta}\log p(y,\theta)$ $=-a_{1}+b_{1}\exp(-\theta).$
Taking the expectation of $\log p(y,\theta)$ with respect to $q(\theta)$,
${\text{E}}_{q}\left\{\log p(y,\theta)\right\}$ $=-\frac{n}{2}\log(2\pi)+a_{0}\log b_{0}-\log\Gamma(a_{0})$
$\quad-a_{1}\mu-\exp\left(\frac{\sigma^{2}}{2}-\mu\right)b_{1}.$
Hence, the evidence lower bound for KLD is
${\mathcal{L}}$ $={\text{E}}_{q}\left\{\log p(y,\theta)-\log q(\theta)\right\}$
$=-\frac{n}{2}\log(2\pi)+a_{0}\log b_{0}-\log\Gamma(a_{0})-a_{1}\mu$
$\quad-\exp\left(\frac{\sigma^{2}}{2}-\mu\right)b_{1}+\frac{1}{2}\log(2\pi\sigma^{2})+\frac{1}{2}.$
For the Fisher divergence, we evaluate the expectation of the squared norm of the gradient difference,
$F(q\|p)$ $={\text{E}}_{q}\{\|\nabla_{\theta}\log p(y,\theta)-\nabla_{\theta}\log q(\theta)\|^{2}\}$
$={\text{E}}_{q}\left\{\left\|-a_{1}+\exp(-\theta)b_{1}+\frac{\theta-\mu}{\sigma^{2}}\right\|^{2}\right\}$
$=a_{1}^{2}+b_{1}^{2}{\text{E}}_{q}\left\{\exp(-2\theta)\right\}+{\text{E}}_{q}\left\{\frac{(\theta-\mu)^{2}}{\sigma^{4}}\right\}$
$\quad-2a_{1}b_{1}{\text{E}}_{q}\left\{\exp(-\theta)\right\}-2a_{1}{\text{E}}_{q}\left\{\frac{\theta-\mu}{\sigma^{2}}\right\}$
$\quad+2b_{1}{\text{E}}_{q}\left\{\exp(-\theta)\frac{\theta-\mu}{\sigma^{2}}\right\}.$
Note that ${\text{E}}_{q}\{\exp(-a\theta)\}=\exp(a^{2}\sigma^{2}/2-a\mu)$ for any constant $a\in\mathbb{R}$. Thus,
$F(q\|p)$ $=a_{1}^{2}+b_{1}^{2}\exp(2\sigma^{2}-2\mu)$
$\quad-2b_{1}(a_{1}+1)\exp(\sigma^{2}/2-\mu)+1/\sigma^{2}.$
We have
$S(q\|p)$ $=\sigma^{2}F(q\|p)$
$=\sigma^{2}\{a_{1}^{2}+b_{1}^{2}\exp(2\sigma^{2}-2\mu)$
$\quad-2b_{1}(a_{1}+1)\exp(\sigma^{2}/2-\mu)\}+1.$
$\nabla_{\mu}S(q\|p)$ $=2b_{1}\sigma^{2}\exp(\sigma^{2}/2-\mu)\{(a_{1}+1)$
$\quad-b_{1}\exp(3\sigma^{2}/2-\mu)\}=0$
$\therefore\mu$ $=\log\frac{b_{1}}{a_{1}+1}+\frac{3\sigma^{2}}{2}.$
At this value of $\mu$,
$S(q\|p)$ $=a_{1}^{2}\sigma^{2}-(a_{1}+1)^{2}\sigma^{2}\exp(-\sigma^{2})+1.$
$\nabla_{\sigma^{2}}S(q\|p)$ $=a_{1}^{2}+(a_{1}+1)^{2}\exp(-\sigma^{2})(\sigma^{2}-1).$
S1.3 Skew normal
The probability density function (pdf) for $\theta\sim{\text{SN}}(m,t,\lambda)$ is
$p(\theta,y)=2\,\phi(\theta|m,t^{2})\,\Phi\{\lambda(\theta-m)\}.$
The log-density and gradient for the skew normal are
$\log p(\theta,y)$ $=\log 2-\frac{1}{2}\log(2\pi t^{2})-\frac{(\theta-m)^{2}}{2t^{2}}$
$\quad+\log[\Phi\{\lambda(\theta-m)\}],$
$\nabla_{\theta}\log p(\theta,y)$ $=-\frac{\theta-m}{t^{2}}+\frac{\lambda\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}.$
Taking the expectation of $\log p(\theta,y)$ with respect to $q(\theta)$,
${\text{E}}_{q}\left\{\log p(\theta,y)\right\}$ $=\log 2-\frac{1}{2}\log(2\pi t^{2})-\frac{\sigma^{2}+(\mu-m)^{2}}{2t^{2}}$
$\quad+{\text{E}}_{q}\log[\Phi\{\lambda(\theta-m)\}].$
For KLD, we maximize the evidence lower bound
${\mathcal{L}}$ $={\text{E}}_{q}\left\{\log p(\theta,y)-\log q(\theta)\right\}$
$=\log 2-\frac{1}{2}\log(2\pi t^{2})-\frac{\sigma^{2}+(\mu-m)^{2}}{2t^{2}}$
$\quad+{\text{E}}_{q}\log[\Phi\{\lambda(\theta-m)\}]+\frac{1}{2}\log(2\pi\sigma^{2})+\frac{1}{2}.$
The Fisher divergence is given by
$F(q\|p)={\text{E}}_{q}\{\|\nabla_{\theta}\log p(\theta,y)-\nabla_{\theta}\log q(\theta)\|^{2}\}$
$={\text{E}}_{q}\left\{\left\|-\frac{\theta-m}{t^{2}}+\frac{\lambda\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}+\frac{\theta-\mu}{\sigma^{2}}\right\|^{2}\right\}$
$={\text{E}}_{q}\left\{\frac{(\theta-m)^{2}}{t^{4}}\right\}+2\,{\text{E}}_{q}\left[\frac{\lambda\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}\frac{(\theta-\mu)}{\sigma^{2}}\right]$
$\quad+{\text{E}}_{q}\left\{\frac{(\theta-\mu)^{2}}{\sigma^{4}}\right\}-2\,{\text{E}}_{q}\left[\frac{(\theta-m)}{t^{2}}\frac{\lambda\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}\right]$
$\quad-2\,{\text{E}}_{q}\left\{\frac{(\theta-m)(\theta-\mu)}{t^{2}\sigma^{2}}\right\}+{\text{E}}_{q}\left[\frac{\lambda^{2}\phi^{2}\{\lambda(\theta-m)\}}{\Phi^{2}\{\lambda(\theta-m)\}}\right].$
After computing the 1st, 3rd and 5th terms in the final expression exactly, we obtain
$F(q\|p)$ $=\frac{\sigma^{2}+(\mu-m)^{2}}{t^{4}}+\lambda^{2}{\text{E}}_{q}\left[\frac{\phi^{2}\{\lambda(\theta-m)\}}{\Phi^{2}\{\lambda(\theta-m)\}}\right]$
$\quad+\frac{1}{\sigma^{2}}-\frac{2\lambda}{t^{2}}\,{\text{E}}_{q}\left[(\theta-m)\frac{\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}\right]-\frac{2}{t^{2}}$
$\quad+\frac{2\lambda}{\sigma^{2}}{\text{E}}_{q}\left[(\theta-\mu)\frac{\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}\right].$
S2 SGD based on reparametrization trick
As $\theta=\mu+T^{-\top}z$, we have
$d\theta=d\mu\end{aligned}\quad\text{and}\quad\begin{aligned} d\theta=-T^{-\top}(dT^{\top})T^{-\top}z.\end{aligned}$
Recall that
$g(\lambda,\theta)$ $=\nabla_{\theta}\log h(\theta)+TT^{\top}(\theta-\mu),$
$f(\lambda,\theta)$ $=T^{-1}\nabla_{\theta}\log h(\theta)+T^{\top}(\theta-\mu).$
Let $\text{vec}(\cdot)$ be the operator that stacks all elements of a matrix into a vector columnwise from left to right. In addition, let $K$ be the commutation matrix such that $K\text{\rm vec}(A)=\text{\rm vec}(A^{\top})$, and $L$ be the elimination matrix such that $L\text{\rm vec}(A)=\text{\rm vech}(A)$ for any $d\times d$ matrix $A$, and $L^{\top}\text{\rm vech}(A)=\text{\rm vec}(A)$ if $A$ is lower triangular.
Differentiating $g(\lambda,\theta)$ w.r.t. $\mu$,
$dg(\lambda,\theta)$ =$=\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}d\theta+TT^{\top}(d\theta-d\mu)$
=$=\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}d\mu$
$\therefore\nabla_{\mu}g(\lambda,\theta)$ $=\nabla^{2}_{\theta}\log h(\theta).$
Differentiating $g(\lambda,\theta)$ w.r.t. $\text{\rm vech}(T)$,
$dg(\lambda,\theta)$ =$=\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}d\theta+(dT)T^{\top}(\theta-\mu)$
$\quad+T(dT^{\top})(\theta-\mu)+TT^{\top}(d\theta)$
=$=-\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}T^{-\top}(dT^{\top})T^{-\top}z$
$\quad+(dT)T^{\top}(\theta-\mu)+T(dT^{\top})(\theta-\mu)$
$\quad-T(dT^{\top})T^{-\top}z$
=$=-\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}T^{-\top}(dT^{\top})T^{-\top}z+(dT)z$
=$=\{-(z^{\top}T^{-1}\otimes\nabla^{2}_{\theta}\log h(\theta)^{\top}T^{-\top})K$
$\quad+(z^{\top}\otimes I_{d})\}L^{\top}d\text{\rm vech}(T)$
=$=\{-(T^{-1}\nabla^{2}_{\theta}\log h(\theta)\otimes T^{-\top}z)$
$\quad+(z\otimes I_{d})\}^{\top}L^{\top}d\text{\rm vech}(T).$
$\therefore\nabla_{\text{\rm vech}(T)}g(\lambda,\theta)$ $=L\{(z\otimes I_{d})$
$\quad-(T^{-1}\nabla^{2}_{\theta}\log h(\theta)\otimes T^{-\top}z)\}.$
Differentiating $f(\lambda,\theta)$ w.r.t. $\mu$,
$df(\lambda,\theta)$ $=T^{-1}\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}d\theta+T^{\top}(d\theta-d\mu)$
$=T^{-1}\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}d\mu.$
$\therefore\nabla_{\mu}f(\lambda,\theta)$ $=\nabla^{2}_{\theta}\log h(\theta)T^{-\top}.$
Differentiating $f(\lambda,\theta)$ w.r.t. $\text{\rm vech}(T)$,
$df(\lambda,\theta)$ $=-T^{-1}(dT)T^{-1}\nabla_{\theta}\log h(\theta)+T^{\top}d\theta$
$\quad+(dT^{\top})(\theta-\mu)+T^{-1}\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}d\theta$
$=-T^{-1}(dT)T^{-1}\nabla_{\theta}\log h(\theta)+(dT^{\top})(\theta-\mu)$
$\quad-T^{-1}\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}T^{-\top}dT^{\top}T^{-\top}z$
$\quad-(dT^{\top})T^{-\top}z$
=$=-\{(z^{\top}T^{-1}\otimes T^{-1}\nabla^{2}_{\theta}\log h(\theta)^{\top}T^{-\top})K$
$\quad+(\nabla_{\theta}\log h(\theta)^{\top}T^{-\top}\otimes T^{-1})\}L^{\top}d\text{\rm vech}(T)$
=$=-\{(T^{-1}\nabla^{2}_{\theta}\log h(\theta)T^{-\top}\otimes T^{-\top}z)$
$\quad+(T^{-1}\nabla_{\theta}\log h(\theta)\otimes T^{-\top})\}^{\top}L^{\top}d\text{\rm vech}(T).$
$\therefore\nabla_{\text{\rm vech}(T)}f(\lambda,\theta)$ $=-L\{(T^{-1}\nabla^{2}_{\theta}\log h(\theta)T^{-\top}\otimes T^{-\top}z)$
$\quad+(T^{-1}\nabla_{\theta}\log h(\theta)\otimes T^{-\top})\}.$
Differentiating
$F(\lambda)={\text{E}}_{\phi}\left\{g(\lambda,\mu+T^{-\top}z)^{\top}g(\lambda,\mu+T^{-\top}z)\right\}$
with respect to $\mu$,
$dF(\lambda)$ $={\text{E}}_{\phi}\left[2g(\lambda,\theta)^{\top}dg(\lambda,\theta)\right]$
$={\text{E}}_{\phi}\left[2g(\lambda,\theta)^{\top}\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}d\mu\right].$
$\therefore\nabla_{\mu}F(\lambda)$ $=2{\text{E}}_{\phi}\left[\{\nabla^{2}_{\theta}\log h(\theta)\}g(\lambda,\theta)\right].$
Next we differentiate $F(\lambda)$ with respect to $\text{\rm vech}(T)$.
$dF(\lambda)$ $={\text{E}}_{\phi}\left[2g(\lambda,\theta)^{\top}dg(\lambda,\theta)\right]$
$=2{\text{E}}_{\phi}\left[g(\lambda,\theta)^{\top}\{-(T^{-1}\nabla^{2}_{\theta}\log h(\theta)\otimes T^{-\top}z)\right.$
$\quad\left.+(z\otimes I_{d})\}^{\top}L^{\top}d\text{\rm vech}(T)\right],$
$\nabla_{\text{\rm vech}(T)}$ $F(\lambda)=2L{\text{E}}_{\phi}\left[\{-(T^{-1}\nabla^{2}_{\theta}\log h(\theta)\otimes T^{-\top}z)\right.$
$\quad\left.+(z\otimes I_{d})\}g(\lambda,\theta)\right]$
$=2{\text{E}}_{\phi}\text{\rm vech}\{-T^{-\top}zg(\lambda,\theta)^{\top}\nabla_{\theta}^{2}\log h(\theta)T^{-\top}$
$\quad+g(\lambda,\theta)z^{\top}\}.$
Differentiating $S(\lambda)$ with respect to $\mu$,
$dS(\lambda)$ $={\text{E}}_{\phi}\left[\{2f(\lambda,\theta)\}^{\top}df(\lambda,\theta)\right]$
$={\text{E}}_{\phi}\left[\{2f(\lambda,\theta)\}^{\top}T^{-1}\{\nabla^{2}_{\theta}\log h(\theta)\}^{\top}d\mu\right],$
$\therefore\nabla_{\mu}S(\lambda)$ $=2{\text{E}}_{\phi}\{\nabla_{\theta}^{2}\log h(\theta)T^{-T}f(\lambda,\theta)\}.$
Differentiating $S(\lambda)$ with respect to $\text{\rm vech}(T)$,
$d$ $S(\lambda)={\text{E}}_{\phi}\left[2f(\lambda,\theta)^{\top}df(\lambda,\theta)\right]$
$=-2{\text{E}}_{\phi}[f(\lambda,\theta)^{\top}\{(T^{-1}\nabla^{2}_{\theta}\log h(\theta)T^{-\top}\otimes T^{-\top}z)$
$\quad+(T^{-1}\nabla_{\theta}\log h(\theta)\otimes T^{-\top})\}^{\top}L^{\top}d\text{\rm vech}(T)],$
$\nabla_{\text{\rm vech}(T)}S(\lambda)$
$=-2L{\text{E}}_{\phi}\left[\{(T^{-1}\nabla^{2}_{\theta}\log h(\theta)T^{-\top}\otimes T^{-\top}z)\right.$
$\quad\left.+(T^{-1}\nabla_{\theta}\log h(\theta)\otimes T^{-\top})\}f(\lambda,\theta)\right]$
$=-2{\text{E}}_{\phi}\text{\rm vech}\{T^{-\top}f(\lambda,\theta)\nabla_{\theta}\log h(\theta)^{\top}T^{-\top}$
$\quad+T^{-\top}zf(\lambda,\theta)^{\top}T^{-1}\nabla_{\theta}^{2}\log h(\theta)T^{-\top}\}.$
S2.1 Variance of gradient estimates
We have
$g_{\mu}^{\text{\rm KL}}$ $=\nabla_{\theta}\log h(\theta)+Tz=-\Lambda(\theta-\nu)+Tz$
$=-\Lambda(T^{-\top}z+\mu-\nu)+Tz$
$=(T-\Lambda T^{-\top})z-\Lambda(\mu-\nu),$
$g_{\mu}^{\text{F}}$ $=-2\nabla_{\theta}^{2}\log h(\theta)g_{\mu}^{\text{\rm KL}}=-2(-\Lambda)g_{\mu}^{\text{\rm KL}}=2\Lambda g_{\mu}^{\text{\rm KL}},$
$g_{\mu}^{\text{S}}$ $=-2\nabla_{\theta}^{2}\log h(\theta)T^{-\top}T^{-1}g_{\mu}^{\text{\rm KL}}$
$=2\Lambda T^{-\top}T^{-1}g_{\mu}^{\text{\rm KL}},$
$g_{T}^{\text{\rm KL}}$ $=T^{-\top}z(\mu-\nu)^{\top}\Lambda T^{-\top}$
$\quad-T^{-\top}zz^{\top}(T^{\top}-T^{-1}\Lambda)T^{-\top},$
$g_{T}^{\text{F}}$ $=2\{\Lambda(\mu-\nu)z^{\top}+T^{-\top}z(\mu-\nu)^{\top}\Lambda^{2}T^{-\top}$
$\quad-(T-\Lambda T^{-\top})zz^{\top}$
$\quad-T^{-\top}zz^{\top}(T^{\top}-T^{-1}\Lambda)\Lambda T^{-\top}\},$
$g_{T}^{\text{S}}$ $=2[-\Sigma(T-\Lambda T^{-\top})\{zz^{\top}T^{-1}+z(\mu-\nu)^{\top}\}$
×$\times\Lambda T^{-\top}+\Sigma\Lambda(\mu-\nu)\{z^{\top}T^{-1}+(\mu-\nu)^{\top}\}\Lambda T^{-\top}$
$+T^{-\top}\{z(\mu-\nu)^{\top}\Lambda-zz^{\top}(T^{\top}-T^{-1}\Lambda)\}\Sigma\Lambda T^{-\top}],$
and
$\text{\rm Var}(g_{\mu}^{\text{\rm KL}})$ $=(T-\Lambda T^{-\top})(T^{\top}-T^{-1}\Lambda)$
$=\Sigma^{-1}-2\Lambda+\Lambda\Sigma\Lambda,$
$\text{\rm Var}(g_{\mu}^{\text{F}})$ $=4\Lambda\text{\rm Var}(g_{\mu}^{\text{\rm KL}})\Lambda,$
$\text{\rm Var}(g_{\mu}^{\text{S}})$ $=4\Lambda T^{-\top}T^{-1}\text{\rm Var}(g_{\mu}^{\text{\rm KL}})T^{-\top}T^{-1}\Lambda.$
To simplify the derivation of the variance with respect to $T_{ii}$, we further assume that both $\Lambda$ and $T$ are diagonal matrices. Under this assumption, the gradient terms can be expressed as
$g_{T_{ii}}^{\text{\rm KL}}$ $=\frac{\Lambda_{ii}(\mu_{i}-\nu_{i})}{T_{ii}^{2}}z_{i}+\left(-\frac{1}{T_{ii}}+\frac{\Lambda_{ii}}{T_{ii}^{3}}\right)z_{i}^{2},$
$g_{T_{ii}}^{\text{F}}$ $=2\left(\Lambda_{ii}+\frac{\Lambda_{ii}^{2}}{T_{ii}^{2}}\right)(\mu_{i}-\nu_{i})z_{i}+2\left(-T_{ii}+\frac{\Lambda_{ii}^{2}}{T_{ii}^{3}}\right)z_{i}^{2}$
$=2(T_{ii}^{2}+\Lambda_{ii})g_{T_{ii}}^{\text{\rm KL}},$
$g_{T_{ii}}^{\text{S}}$ $=2\frac{\Lambda_{ii}^{2}(\mu_{i}-\nu_{i})^{2}}{T_{ii}^{3}}+2z_{i}(\mu_{i}-\nu_{i})\left(-\frac{\Lambda_{ii}}{T_{ii}^{2}}+3\frac{\Lambda_{ii}^{2}}{T_{ii}^{4}}\right)$
$\quad+4z_{i}^{2}\left(-\frac{\Lambda_{ii}}{T_{ii}^{3}}+\frac{\Lambda_{ii}^{2}}{T_{ii}^{5}}\right).$
Utilizing the properties $\text{\rm Var}(z_{i})=1$, $\text{\rm Var}(z_{i}^{2})=2$ and $\text{\rm cov}(z_{i},z_{i}^{2})=0$, we obtain
$\text{\rm Var}(g_{T_{ii}}^{\text{\rm KL}})$ $=\frac{1}{T_{ii}^{4}}\left\{\Lambda_{ii}^{2}(\mu_{i}-\nu_{i})^{2}+2\left(T_{ii}-\frac{\Lambda_{ii}}{T_{ii}}\right)^{2}\right\},$
$\text{\rm Var}(g_{T_{ii}}^{\text{F}})$ $=4(T_{ii}^{2}+\Lambda_{ii})^{2}\text{\rm Var}(g_{T_{ii}}^{\text{\rm KL}}),$
$\text{\rm Var}(g_{T_{ii}}^{\text{S}})$ $=\frac{4\Lambda_{ii}^{2}}{T_{ii}^{8}}\left\{\left(3\Lambda_{ii}-T_{ii}^{2}\right)^{2}(\mu_{i}-\nu_{i})^{2}\right.$
$\quad\left.+8\left(T_{ii}-\frac{\Lambda_{ii}}{T_{ii}}\right)^{2}\right\}.$
S3 SGD based on batch approximation
We have
$\hat{S}_{q_{t}}(\lambda)=\frac{1}{B}\sum_{b=1}^{B}\{g_{h}(\theta_{i})^{\top}\Sigma g_{h}(\theta_{i})+2g_{h}(\theta_{i})^{\top}(\theta_{i}-\mu)$
$\quad+(\theta_{i}-\mu)^{\top}\Sigma^{-1}(\theta_{i}-\mu)\}$
$=\frac{1}{B}\sum_{b=1}^{B}[\{g_{h}(\theta_{i})-\overline{g}_{h}+\overline{g}_{h}\}^{\top}\Sigma\{g_{h}(\theta_{i})-\overline{g}_{h}+\overline{g}_{h}\}$
$\quad+2\{g_{h}(\theta_{i})-\overline{g}_{h}+\overline{g}_{h}\}^{\top}(\theta_{i}-\overline{\theta}+\overline{\theta}-\mu)$
$\quad+(\theta_{i}-\overline{\theta}+\overline{\theta}-\mu)^{\top}\Sigma^{-1}(\theta_{i}-\overline{\theta}+\overline{\theta}-\mu)]$
$=\text{\rm tr}\{(C_{g}+\overline{g}_{h}\overline{g}_{h}^{\top})\Sigma\}+\text{\rm tr}(C_{\theta}\Sigma^{-1})$
$\quad+(\mu-\overline{\theta})^{\top}\Sigma^{-1}(\mu-\overline{\theta})-2\overline{g}_{h}^{\top}(\mu-\overline{\theta})$
$\quad+\frac{2}{B}\sum_{b=1}^{B}\{g_{h}(\theta_{i})-\overline{g}_{h}\}^{\top}(\theta_{i}-\overline{\theta})$
$=\text{\rm tr}(V\Sigma)+\text{\rm tr}(U\Sigma^{-1})+2\text{\rm tr}(W),$
$\hat{F}_{q_{t}}(\lambda)=\frac{1}{B}\sum_{b=1}^{B}\{2g_{h}(\theta_{i})^{\top}\Sigma^{-1}(\theta_{i}-\mu)$
$\quad+(\theta_{i}-\mu)^{\top}\Sigma^{-2}(\theta_{i}-\mu)\}+g_{h}(\theta_{i})^{\top}g_{h}(\theta_{i})$
$=\frac{1}{B}\sum_{b=1}^{B}[\{g_{h}(\theta_{i})-\overline{g}_{h}+\overline{g}_{h}\}^{\top}\{g_{h}(\theta_{i})-\overline{g}_{h}+\overline{g}_{h}\}$
$\quad+2\{g_{h}(\theta_{i})-\overline{g}_{h}+\overline{g}_{h}\}^{\top}\Sigma^{-1}(\theta_{i}-\overline{\theta}+\overline{\theta}-\mu)$
$\quad+(\theta_{i}-\overline{\theta}+\overline{\theta}-\mu)^{\top}\Sigma^{-2}(\theta_{i}-\overline{\theta}+\overline{\theta}-\mu)]$
$=\text{\rm tr}(C_{g}+\overline{g}_{h}\overline{g}_{h}^{\top})+2\text{\rm tr}(C_{\theta g}\Sigma^{-1})+\text{\rm tr}(C_{\theta}\Sigma^{-2})$
$\quad+(\mu-\overline{\theta})^{\top}\Sigma^{-2}(\mu-\overline{\theta})-2\overline{g}_{h}^{\top}\Sigma^{-1}(\mu-\overline{\theta})$
$=\text{\rm tr}(V)+\text{\rm tr}(U\Sigma^{-2})+2\text{\rm tr}(W\Sigma^{-1}),$
where $U=C_{\theta}+(\mu-\overline{\theta})(\mu-\overline{\theta})^{\top}$, $V=C_{g}+\overline{g}_{h}\overline{g}_{h}^{\top}$ and $W=C_{\theta g}-(\mu-\overline{\theta})\overline{g}_{h}^{T}$. Note that $U$ and $V$ are symmetric but $W$ is not. Differentiating with respect to $\mu$ and $T$,
$\nabla_{\mu}\hat{S}_{q_{t}}(\lambda)$ $=2\Sigma^{-1}(\mu-\overline{\theta})-2\overline{g}_{h}.$
$d\hat{S}_{q_{t}}(\lambda)$ $=d\{\text{\rm tr}(VT^{-\top}T^{-1})+\text{\rm tr}(UTT^{\top})\}$
$=-\text{\rm tr}(VT^{-\top}dT^{\top}\Sigma)-\text{\rm tr}(V\Sigma dTT^{-1})$
$\quad+\text{\rm tr}(UdTT^{\top})+\text{\rm tr}(UTdT^{\top})$
$=\text{\rm tr}\{(UT-\Sigma VT^{-\top})dT^{\top}\}$
$\quad+\text{\rm tr}\{(T^{\top}U-T^{-1}V\Sigma)dT\}$
$=2\text{\rm vec}(UT-\Sigma VT^{-\top})^{\top}L^{\top}d\text{\rm vech}(T).$
$\nabla_{\text{\rm vech}(T)}$ $\hat{S}_{q_{t}}(\lambda)=2\text{\rm vech}(UT-\Sigma VT^{-\top}).$
$\nabla_{\mu}\hat{F}_{q_{t}}(\lambda)$ $=\Sigma^{-1}\nabla_{\mu}\hat{S}_{q_{t}}(\lambda).$
$d\hat{F}_{q_{t}}(\lambda)$ $=d\{\text{\rm tr}(UTT^{\top}TT^{\top})+2\text{\rm tr}(WTT^{\top})\}$
$=\text{\rm tr}(UdTT^{\top}\Sigma^{-1})+\text{\rm tr}(UTdT^{\top}\Sigma^{-1})$
$\quad+\text{\rm tr}(U\Sigma^{-1}dTT^{\top})+\text{\rm tr}(U\Sigma^{-1}TdT^{\top})$
$\quad+2\text{\rm tr}(WdTT^{\top})+2\text{\rm tr}(WTdT^{\top})$
$=\text{\rm tr}\{(T^{\top}\Sigma^{-1}U+T^{\top}U\Sigma^{-1})dT\}$
$\quad+\text{\rm tr}\{(\Sigma^{-1}UT+U\Sigma^{-1}T)dT^{\top}\}$
$\quad+2\text{\rm tr}(T^{\top}WdT)+2\text{\rm tr}(WTdT^{\top})$
$=2\text{\rm vec}(\Sigma^{-1}UT+U\Sigma^{-1}T+W^{\top}T$
$\quad+WT)^{\top}L^{\top}d\text{\rm vech}(T).$
$\nabla_{\text{\rm vech}(T)}\hat{F}_{q_{t}}(\lambda)$ $=2\text{\rm vech}\{(W+W^{\top}+\Sigma^{-1}U$
$\quad+U\Sigma^{-1})T\}.$
S4 Gradients for GLMMs
Let $X_{i}=(X_{i1},\dots,X_{in_{i}})^{\top}$ and $Z_{i}=(Z_{i1},\dots,Z_{in_{i}})^{\top}$ be design matrices for the $i$th subject. Recall that $b_{i}\sim{\text{N}}(0,G^{-1})$, $G=WW^{\top}$, $W^{*}$ is such that $W_{ii}^{*}=\log(W_{ii})$ and $W_{ij}^{*}=W_{ij}$ if $i\neq j$, and $\zeta=\text{vech}(W^{*})$. Let $J^{W}$ be an $r\times r$ matrix with diagonal given by $\text{\rm diag}(W)$ and all off-diagonal entries being 1, and $D^{W}=\text{\rm diag}\{\text{\rm vech}(J^{W})\}$. Then $d\text{\rm vech}(W)=D^{W}d\text{\rm vech}(W^{*})$. We have
$\nabla_{\theta}\log h(\theta)=[\nabla_{b_{1}}\log h(\theta)^{\top},\dots,\nabla_{b_{n}}\log h(\theta)^{\top},\\ \nabla_{\beta}\log h(\theta)^{\top},\nabla_{\zeta}\log h(\theta)^{\top}]^{\top},$
where
$\nabla_{b_{i}}\log h(\theta)$ $=\sum_{j=1}^{n_{i}}\{y_{ij}-A^{\prime}(\eta_{ij})\}Z_{ij}-Gb_{i},$
$\nabla_{\beta}\log h(\theta)$ $=\sum_{i=1}^{n}\sum_{j=1}^{n_{i}}\{y_{ij}-A^{\prime}(\eta_{ij})\}X_{ij}-\frac{\beta}{\sigma_{\beta}^{2}},$
$\nabla_{\zeta}\log h(\theta)$ $=-D^{W}\text{\rm vech}(\widetilde{W})+n\text{\rm vech}(I_{r})-\frac{\zeta}{\sigma_{\zeta}^{2}},$
and $\widetilde{W}=\sum_{i=1}^{n}b_{i}b_{i}^{\top}W$. For the hessian $\nabla_{\theta}^{2}\log h(\theta)$,
$\nabla_{b_{i},\eta}^{2}\log h(\theta)$ $=\begin{bmatrix}\nabla_{b_{i},\beta}^{2}\log h(\theta)\\ \nabla_{b_{i},\zeta}^{2}\log h(\theta)\\ \end{bmatrix},$
$\nabla_{\eta}^{2}\log h(\theta)$ $=\begin{bmatrix}\nabla_{\beta}^{2}\log h(\theta)&0\\ 0&\nabla_{\zeta}^{2}\log h(\theta)\\ \end{bmatrix}.$
Let $B_{i}=\text{\rm diag}([A^{\prime\prime}(\eta_{i1}),\dots,A^{\prime\prime}(\eta_{in_{i}})]^{\top})$. The second order derivatives of $\log h(\theta)$ are
$\nabla_{b_{i}}^{2}\log h(\theta)$ $=-(Z_{i}^{\top}B_{i}Z_{i}+G),\text{ for },$
$\nabla_{\beta}^{2}\log h(\theta)$ $=-\left(\sum_{i=1}^{n}X_{i}^{\top}B_{i}X_{i}+\frac{1}{\sigma^{2}_{\beta}}I_{p}\right),$
$\nabla_{\zeta}^{2}\log h(\theta)$ $=-S-D^{W}L\sum_{i=1}^{n}(I_{r}\otimes b_{i}b_{i}^{\top})L^{\top}D^{W}$
$\quad-\frac{1}{\sigma^{2}_{\zeta}}I_{r(r+1)/2},$
$\nabla_{\beta,b_{i}}^{2}\log h(\theta)$ $=-X_{i}^{\top}B_{i}Z_{i},$
$\nabla_{\zeta,b_{i}}^{2}\log h(\theta)$ $=-D^{W}L(W^{\top}b_{i}\otimes I_{r}+W^{\top}\otimes b_{i}),$
where $S=\text{\rm diag}[\text{\rm vech}\{\text{\rm dg}(W)\text{\rm dg}(\widetilde{W})\}]$ and $\text{\rm dg}(A)$ is a copy of $A$ with all off-diagonal entries set to 0.
The derivations for $\nabla_{b_{i}}^{2}\log h(\theta)$, $\nabla_{\beta}^{2}\log h(\theta)$ and $\nabla_{b_{i},\beta}^{2}\log h(\theta)$ are straightforward. More details for $\nabla_{\zeta^{2}}^{2}\log h(\theta)$ and $\nabla_{b_{i},\zeta}^{2}\log h(\theta)$ are given below. Differentiating $\nabla_{\zeta}\log h(\theta)$ w.r.t. $b_{i}$, we have
$d\nabla_{\zeta}$ $\log h(\theta)=-D^{W}\sum_{i=1}^{n}\text{\rm vech}\{(db_{i})b_{i}^{\top}W+b_{i}(db_{i}^{\top})W\}$
$=-D^{W}L\sum_{i=1}^{n}[(W^{\top}b_{i}\otimes I_{r})+(W^{\top}\otimes b_{i})]db_{i}.$
Differentiating $\nabla_{\zeta}\log h(\theta)$ w.r.t. $\zeta$, we have
$d\nabla_{\zeta}\log h(\theta)=-(dD^{W})\sum_{i=1}^{n}\text{\rm vech}(b_{i}b_{i}^{\top}W)$
$\quad-D^{W}\sum_{i=1}^{n}\text{\rm vech}\{b_{i}b_{i}^{\top}(dW)\}-\frac{1}{\sigma^{2}_{\zeta}}d\zeta$
$=-D^{W}L\sum_{i=1}^{n}(I_{r}\otimes b_{i}b_{i}^{\top})d\text{\rm vec}(W)$
$\quad-Sd\zeta-\frac{1}{\sigma^{2}_{\zeta}}d\zeta$
$=-D^{W}L\sum_{i=1}^{n}(I_{r}\otimes b_{i}b_{i}^{\top})L^{T}D^{W}d\zeta$
$\quad-Sd\zeta-\frac{1}{\sigma^{2}_{\zeta}}d\zeta.$
S5 Gradients for stochastic volatility model
For this model, the gradients of $\log h(\theta)$ are,
$\nabla_{b_{1}}\log h(\theta)$ $=-(1-\phi^{2})b_{1}+\phi(b_{2}-\phi b_{1})-\frac{e^{\alpha}}{2}$
$\quad+\frac{e^{\alpha}y_{1}^{2}}{2}\exp(-\lambda-e^{\alpha}b_{1}),$
$\nabla_{b_{t}}\log h(\theta)$ $=\phi(b_{t+1}-\phi b_{t})-(b_{t}-\phi b_{t-1})-\frac{e^{\alpha}}{2}$
$\quad+\frac{e^{\alpha}}{2}y_{t}^{2}\exp(-\lambda-e^{\alpha}b_{t})\;\;\text{for}\;1<t<n,$
$\nabla_{b_{n}}\log h(\theta)$ $=-(b_{n}-\phi b_{n-1})-\frac{e^{\alpha}}{2}$
$\quad+\frac{e^{\alpha}}{2}y_{n}^{2}\exp(-\lambda-e^{\alpha}b_{n}),$
$\nabla_{\alpha}\log h(\theta)$ $=\frac{1}{2}\sum_{t=1}^{n}y_{t}^{2}b_{t}\exp(\alpha-\lambda-e^{\alpha}b_{t})$
$\quad-\frac{e^{\alpha}}{2}\sum_{t=1}^{n}b_{t}-\frac{\alpha}{\sigma_{0}^{2}},$
$\nabla_{\lambda}\log h(\theta)$ $=-\frac{n}{2}+\frac{1}{2}\sum_{t=1}^{n}y_{t}^{2}\exp(-\lambda-e^{\alpha}b_{t})-\frac{\lambda}{\sigma_{0}^{2}},$
$\nabla_{\psi}\log h(\theta)$ =$=\Big{\{}\phi b_{1}^{2}-\frac{\phi}{(1-\phi^{2})}+\sum_{t=1}^{n-1}(b_{t+1}-\phi b_{t})b_{t}\Big{\}}$
×$\quad\times\frac{e^{\psi}}{(e^{\psi}+1)^{2}}-\frac{\psi}{\sigma_{0}^{2}}.$
The second order derivatives of $\log h(\theta)$ are,
$\nabla_{b_{1}}^{2}\log h(\theta)$ $=-1-\frac{y_{1}^{2}}{2}\exp\{2\alpha-\lambda-e^{\alpha}b_{1}\},$
$\nabla_{b_{t}}^{2}\log h(\theta)$ $=-\phi^{2}-1-y_{t}^{2}\exp\{2\alpha-\lambda-e^{\alpha}b_{t}\}/2,$
$\nabla_{b_{n}}^{2}\log h(\theta)$ $=-1-y_{n}^{2}\exp\{2\alpha-\lambda-e^{\alpha}b_{n}\}/2,$
$\nabla_{b_{i},b_{j}}^{2}\log h(\theta)$ $=\phi\mathbbm{1}_{|i-j|=1},$
$\nabla_{b_{t},\alpha}^{2}\log h(\theta)$ $=\frac{y_{t}^{2}}{2}\exp\{\alpha-\lambda-e^{\alpha}b_{t}\}(1-b_{t}e^{\alpha})-\frac{e^{\alpha}}{2},$
$\nabla_{b_{t},\lambda}^{2}\log h(\theta)$ $=-y_{t}^{2}\exp\{\alpha-\lambda-e^{\alpha}b_{t}\}/2$
$\nabla_{b_{1},\psi}^{2}\log h(\theta)$ $=\frac{b_{2}e^{\psi}}{(e^{\psi}+1)^{2}},$
$\nabla_{b_{t},\psi}^{2}\log h(\theta)$ $=\frac{e^{\psi}(b_{t+1}-2\phi b_{t}+b_{t-1})}{(e^{\psi}+1)^{2}},$
$\nabla_{b_{n},\psi}^{2}\log h(\theta)$ $=\frac{e^{\psi}b_{n-1}}{(e^{\psi}+1)^{2}},$
$\nabla_{\alpha}^{2}\log h(\theta)$ $=\frac{1}{2}\sum_{t=1}^{n}y_{t}^{2}b_{t}\exp\{\alpha-\lambda-e^{\alpha}b_{t}\}(1-e^{\alpha}b_{t})$
$\quad-\frac{e^{\alpha}}{2}\sum_{t=1}^{n}b_{t}-\frac{1}{\sigma_{0}^{2}},$
$\nabla_{\lambda}^{2}\log h(\theta)$ $=-\frac{1}{2}\sum_{t=1}^{n}y_{t}^{2}\exp(-\lambda-e^{\alpha}b_{t})-\frac{1}{\sigma^{2}_{0}},$
$\nabla_{\psi}^{2}\log h(\theta)$ =$=\bigg{\{}b_{1}^{2}-\sum_{t=1}^{n-1}b_{t}^{2}-\frac{1+\phi^{2}}{(1-\phi^{2})^{2}}\bigg{\}}\frac{e^{2\psi}}{(e^{\psi}+1)^{4}}$
+$+\bigg{\{}\phi b_{1}^{2}-\frac{\phi}{(1-\phi^{2})}+\sum_{t=1}^{n-1}(b_{t+1}-\phi b_{t})b_{t}\bigg{\}}$
×$\times\frac{e^{\psi}(1-e^{\psi})}{(e^{\psi}+1)^{3}}-\frac{1}{\sigma_{0}^{2}},$
$\nabla_{\alpha,\lambda}^{2}\log h(\theta)$ $=-\frac{1}{2}\sum_{t=1}^{n}y_{t}^{2}b_{t}\exp\{\alpha-\lambda-e^{\alpha}b_{t}\},$
$\nabla_{\psi,\lambda}^{2}\log h(\theta)$ $=\nabla_{\psi,\alpha}^{2}\log h(\theta)=0.$